### v1

# Emulation Capacity between Idempotent Channels

Abstract
We study the optimal rates of emulation (also called interconversion) between quantum channels. When the source and the target channels are idempotent, we give a single-letter expression for the zero-error emulation capacity in terms of structural properties of the range of the two channels. This expression shows that channel emulation is not reversible for general idempotent channels. Furthermore, we establish a strong converse rate that matches with the zero-error emulation capacity when the source or the target channel is either an identity or a completely dephasing channel.
1 Introduction
Given target and source quantum channels $\mathcal{F}$ and $\mathcal{G}$ (respectively), a central problem in quantum information theory is to understand the optimal interconversion rates $k/n$ such that one can use $n$ copies of $\mathcal{G}$, combined with an encoder and a decoder channel, to emulate the action of $k$ copies of $\mathcal{F}$. In the special case where $\mathcal{F}$ is a perfect (noiseless) channel, this corresponds to the problem of channel coding for $\mathcal{G}$, which is central in Shannon theory [13]. Symmetrically, in the special case where $\mathcal{G}$ is the identity channel, this corresponds to the problem of channel simulation for $\mathcal{F}$ which is also well-studied within both classical and quantum Shannon theory [1, 2, 3, 4].
In this article, we study the interconversion problem for more general channels. We start by considering the zero-error regime. We say that a channel $\mathcal{G}$ can emulate another channel $\mathcal{F}$ with rate $r$, if there exist an encoding and a decoding channel $\mathcal{E}$ and $\mathcal{D}$ (respectively) as well as a pair of integers $(k,n)$ such that
$\mathcal{F}^{\otimes k}=\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E},$
and $r\leq k/n$. Such a rate $r$ is called achievable. The supremum of achievable rates is called the zero-error emulation capacity of $\mathcal{F}$ by $\mathcal{G}$ and is denoted $C(\mathcal{G}\mapsto\mathcal{F})$. We note that, unlike most works in the literature on reverse Shannon theory [1], the model we consider here does not allow shared randomness or entanglement between the encoder and decoder. The first contribution of this work is a single-letter expression of $C(\mathcal{G}\mapsto\mathcal{F})$ when both $\mathcal{F}$ and $\mathcal{G}$ are assumed to be idempotent, i.e. they satisfy $\mathcal{F}\circ\mathcal{F}=\mathcal{F}$ and $\mathcal{G}\circ\mathcal{G}=\mathcal{G}$. Note that for an arbitrary channel $\Phi$, the sequential composition $\Phi^{n}=\Phi\circ\cdots\circ\Phi$ has a subsequence that converges as $n\to\infty$ to an idempotent channel [18]. The capacities of sequential composition of channels have been studied in [15, 7, 14].
Theorem 1.
Let $\mathcal{F}=\mathcal{F}\circ\mathcal{F}$ and $\mathcal{G}=\mathcal{G}\circ\mathcal{G}$ be two idempotent channels with $\lambda(\mathcal{F})\neq(1)$ [^1] If $\lambda(\mathcal{F})=(1)$, then $\mathcal{F}$ is a replacer channel (see Proposition 1) and it is simple to see that we can emulate it with any other idempotent channel $\mathcal{G}$ with an arbitrarily large rate., then
$C(\mathcal{G}\mapsto\mathcal{F})=\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}.$ (1)
In particular, when $\mathcal{F}$ or $\mathcal{G}$ is either the identity channel or the completely dephasing channel, it suffices to restrict the infimum to $p\in\{1,+\infty\}$:
$C(\mathcal{G}\mapsto\mathcal{F})=\min_{p\in\{1,+\infty\}}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}.$ (2)
[…]
A remarkable aspect about this result is that it is a single-letter formula, i.e. the expression does not involve a limit. This comes from the fact that the norm of the shape vector is multiplicative under tensor product. Another implication is that the emulation capacity of idempotent channels is additive under tensor products of source channels, i.e. $C(\mathcal{G}_{1}\otimes\mathcal{G}_{2}\mapsto\mathcal{F})=C(\mathcal{G}_{1}\mapsto\mathcal{F})+C(\mathcal{G}_{2}\mapsto\mathcal{F})$.
One interesting consequence of this formula is that emulating $\mathcal{F}$ with $\mathcal{G}$ is in general not reversible. In fact, as shown in Example 1, there exists idempotent channels for which $C(\mathcal{G}\mapsto\mathcal{F})C(\mathcal{F}\mapsto\mathcal{G})^{-1}<1$. This follows from the fact that the infimum in the formulas for $C(\mathcal{G}\mapsto\mathcal{F})$ and $C(\mathcal{F}\mapsto\mathcal{G})$ can be reached for different values of $p$.
Next, we consider the setting beyond the zero-error regime, where the emulation does not need to be exact: the rate $r$ is achievable with error $\delta\in[0,1]$ if there exists channel $\mathcal{E},\mathcal{D}$ and $k,n$ with $r\leq k/n$ such that
$\frac{1}{2}\|\mathcal{F}^{\otimes k}-\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}\|_{\diamond}\leq\delta,$
where $\|\cdot\|_{\diamond}$ is the diamond norm (the completely bounded norm between trace classes). We believe that for idempotent channels, the capacity (with vanishing error) should be the same as the zero-error case. Here, we only proved a weaker statement: the capacity cannot be larger than the same formula except that we only take the minimum over $p\in\{1,+\infty\}$. More specifically, we establish a strong converse rate for the emulation capacity.
Theorem 2 (Strong converse rate).
Let $\mathcal{F}=\mathcal{F}\circ\mathcal{F}$ and $\mathcal{G}=\mathcal{G}\circ\mathcal{G}$ be two idempotent channels with $\lambda(\mathcal{F})\neq(1)$. Let $(k_{\nu})_{\nu\in\mathbb{N}}$ and $(n_{\nu})_{\nu\in\mathbb{N}}$ be two integer sequences with $\lim_{\nu\to\infty}n_{\nu}=+\infty$ such that there exists $\operatorname{\varepsilon}>0$ satisfying:
$\lim_{\nu\to\infty}\frac{k_{\nu}}{n_{\nu}}\geq\min_{p\in\{1,+\infty\}}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}+\operatorname{\varepsilon}.$ (3)
Then, for all sequences of encoding and decoding channels $(\mathcal{E}_{\nu})_{\nu\in\mathbb{N}}$, $(\mathcal{D}_{\nu})_{\nu\in\mathbb{N}}$,
$\frac{1}{2}\lim_{\nu\to\infty}\|\mathcal{F}^{\otimes k_{\nu}}-\mathcal{D}_{\nu}\mathcal{G}^{\otimes n_{\nu}}\mathcal{E}_{\nu}\|_{\diamond}=1.$ (4)
Note that for any two channels $\Phi$ and $\Psi$, we have that $\|\Phi-\Psi\|_{\diamond}\leq 2$. Therefore, Theorem 2 states that if one tries to emulate $\mathcal{F}$ with $\mathcal{G}$ with a rate greater than $\log(\|\lambda(\mathcal{G})\|_{p})/\log(\|\lambda(\mathcal{F})\|_{p})$ for $p\in\{1,+\infty\}$, the error in the limit of large block-length is maximal.
If $\mathcal{F}$ or $\mathcal{G}$ are identity channels or dephasing channels, then the achievability result of Theorem 1 and the strong converse rate in Theorem 2 match. In particular, for the case of $\mathcal{F}=\mathrm{Id}$, we obtain an expression for the quantum capacity of arbitrary idempotent channels together with a strong converse. In the case of self-adjoint idempotent channels (also known as tracial conditional expectation), this was proved in [8] (see also [12]).
We conjecture that Theorem 2 can be improved to take an infimum over all $p\in[1,+\infty]$, which would imply that the zero-error achievability in Theorem 1 is tight for all idempotent channels $\mathcal{F}$, $\mathcal{G}$ even if vanishing error is allowed.
The techniques used in our proofs stem from the theory of operator algebras. Nevertheless, familiarity with this theory is not required to understand the present article. We will introduce in Section 2 the basic theory of operator algebras that we use throughout our work.
This article is structured as follows. In Section 2, we introduce our notations and basic preliminaries on quantum information and operator algebra theories. In Section 3, we discuss the decomposition properties of idempotent channels. In Section 4, we prove Theorem 1. We prove Theorem 2 in Section 5. Finally, we apply the theory of approximate $C^{*}$-algebras recently developed by Kitaev [9] to strengthen the converse of Theorem 1 in the appendix.
2 Preliminaries and notations
All Hilbert spaces in this work are assumed to be finite-dimensional. Given a Hilbert space $\mathcal{H}$, we write $\mathcal{L}(\mathcal{H})$ the set of linear operators on $\mathcal{H}$. This set is endowed with the Hilbert-Schmidt inner product, denoted $\langle x,y\rangle:=\operatorname{\operatorname{tr}}(x^{*}y)$ for $x,y\in\mathcal{L}(\mathcal{H})$, with $x^{*}$ the conjugate transpose of $x$ and $\operatorname{\operatorname{tr}}(\cdot)$ the trace on $\mathcal{L}(\mathcal{H})$. Note that $x^{*}$ is the adjoint of $x$ for the Hermitian product of $\mathcal{H}$. We denote $\mathds{1}\in\mathcal{L}(\mathcal{H})$ the identity operator on $\mathcal{H}$. Given $x\in\mathcal{L}(\mathcal{H})$, we call the (right) support of $x$, denoted $\operatorname{\operatorname{supp}}(x)\subseteq\mathcal{H}$, the orthogonal complement of its (right) kernel. We write $x^{0}$ the orthogonal projector on the support of $x$, parallel to its kernel. Furthermore, for $x\in\mathcal{L}(\mathcal{H})$, we write $\|x\|_{1}$ for its Schatten $1$-norm and $\|x\|$ its Schatten $\infty$-norm, which coincide with its operator norm as a linear operator on $\mathcal{H}$. Recall that the set of Hermitian operators in $\mathcal{L}(\mathcal{H})$ is a partially ordered set: for $x=x^{*},y=y^{*}\in\mathcal{L}(\mathcal{H})$, we write $x\leq y$ whenever yxy-x is positive semidefinite.
A state is represented by a positive semidefinite operator $\rho\in\mathcal{L}(\mathcal{H})$ with trace one, which is called a density operator. We will take $\mathcal{D}(\mathcal{H})\subseteq\mathcal{L}(\mathcal{H})$ to be the set of density operators belonging to $\mathcal{L}(\mathcal{H})$. A measurement on a physical system $\mathcal{L}(\mathcal{H})$ is represented by a Positive Operator Valued Measure, POVM for short, which is a set of positive operators $\mu_{1},\dots,\mu_{K}\in\mathcal{L}(\mathcal{H})$, called effects, which sum to the identity, i.e. $\sum_{k=1}^{K}\mu_{k}=\mathds{1}$.
[…]
$\|\mathcal{F}\|_{\diamond}=\sup_{n\in\mathbb{N}^{*}}\|\mathcal{F}\otimes\mathrm{Id}_{n}\|_{1\to 1},$
[…]
$\|\mathcal{F}\|_{1\to 1}=\sup_{\|x\|_{1}\leq 1}\|\mathcal{F}(x)\|_{1}.$
Then, we recall basic properties of the $\ell_{p}$-norms. For $u\in{\mathbb{C}}^{n}$ and $p\in[1,+\infty)$, the $\ell_{p}$-norm of $u$ is defined as
$\|u\|_{p}=\Bigl(\sum_{k=1}^{n}|u_{k}|^{p}\Bigr)^{\frac{1}{p}},$ (5)
and $\|u\|_{\infty}=\sup_{k}|u_{k}|$. In the discussions of the next sections, we will frequently make use of the fact that the $\ell_{p}$-norm are multiplicative under the tensor product of vectors. Given $u\in{\mathbb{C}}^{n}$, $v\in{\mathbb{C}}^{m}$, the tensor product $u\otimes v\in{\mathbb{C}}^{nm}$ is the vector whose coordinates are the products of the coordinates of $u$ and $v$, i.e. $(u_{k}v_{l})_{k,l}$ for $k\in[n]$, $l\in[m]$. Then, for all $p\in[1,+\infty]$, $\|u\otimes v\|_{p}=\|u\|_{p}\|v\|_{p}$.
Finally, we end this section by introducing the notions of finite-dimensional *-algebras and their morphisms.
A *-algebra $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H})$ is a $\mathbb{C}$-linear subspace of operators which is stable under the multiplication and and the * operation of $\mathcal{L}(\mathcal{H})$, i.e. taking the transpose conjugate. That is, for all $x,y\in\mathcal{A}$, $xy\in\mathcal{A}$ and for all $x\in\mathcal{A}$, $x^{*}\in\mathcal{A}$. Furthermore, all *-algebras in this work are assumed to be unital, that is there is an identity $\mathds{1}\in\mathcal{A}$ satisfying $\mathds{1}^{*}=\mathds{1}$ and, for all $x\in\mathcal{A}$, $\mathds{1}x=x\mathds{1}=x$. Such an identity operator is necessarily unique in each *-algebra.
Given two *-algebras $\mathcal{A}$ and $\mathcal{B}$, a *-homomorphism $\iota:\mathcal{A}\rightarrow\mathcal{B}$ is a linear map that preserves the product and the * operation in $\mathcal{A}$. That is, for all $x,y\in\mathcal{A}$, $\iota(xy)=\iota(x)\iota(y)$ and $\iota(x^{*})=\iota(x)^{*}$. In many texts, *-homomorphisms between unital *-algebras are assumed to be unital as well, that is they are assumed to map the identity of $\mathcal{A}$ on the identity of $\mathcal{B}$. Nevertheless, in this article, we will work with more general subunital *-homomorphisms. Such a map $\iota:\mathcal{A}\rightarrow\mathcal{B}$ maps the identity of $\mathcal{A}$ on $\iota(\mathds{1}_{\mathcal{A}})$, which satisfies $\iota(\mathds{1}_{\mathcal{A}})\leq\mathds{1}_{\mathcal{B}}$. Note that *-homomorphisms are completely positive [5].
We summarized all the notations we use throughout this article in Table 1.
$\mathcal{H}$ A finite-dimensional Hilbert space
$\mathcal{L}(\mathcal{H})$ Set of linear operators on $\mathcal{H}$
$\mathcal{D}(\mathcal{H})$ Set of density operators in $\mathcal{L}(\mathcal{H})$
$\mathcal{A},\mathcal{B}$ *-algebras
$\rho$, $\sigma$ Density operators
$\operatorname{\operatorname{supp}}(x)$ Support of $x$
$x^{0}$ Orthogonal projector on the support of $x$
$x^{*}$ Hermitian conjugate of $x$
$\|\cdot\|_{p}$ $\ell_{p}$-norm Ref. eq. (5)
$\mathcal{F}^{*}$ Adjoint of $\mathcal{F}$ for the Hilbert-Schmidt product
$\operatorname{Rg}(\mathcal{F})$ Range of $\mathcal{F}$ Ref. eq. (6)
$\lambda(\mathcal{F})$ Shape vector of $\mathcal{F}$ Ref. Prop. 1
$[n]$ The set of integers $\{1,\dots,n\}$
Table 1: Notations
3 Decomposition properties of idempotent channels
In this section, we discuss the decomposition properties of the range of idempotent channels, which are summarized in the following propositions. The range of a map $f:\Omega\rightarrow\Omega^{\prime}$, denoted $\operatorname{Rg}(f)$, is the set of all possible images over the domain $\Omega$:
$\operatorname{Rg}(f)=\{f(x)~|~x\in\Omega\}.$ (6)
Proposition 1 (Structure of the range of idempotent channels).
[…]
$\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$ (7)
and a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$, such that for all $x\in\mathcal{L}(\mathcal{H})$:
$\mathcal{F}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}},$ (8)
and
$\mathcal{F}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k},$ (9)
where, for all $k\in[K]$, $P_{k}$ is the orthogonal projector onto $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, $d_{k}=\dim(\mathcal{H}_{k,1})\geq 1$, $m_{k}=\dim(\mathcal{H}_{k,2})$ and $\operatorname{\operatorname{tr}}_{k,2}:\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})\rightarrow\mathcal{L}(\mathcal{H}_{k,1})$ is the partial trace over $\mathcal{H}_{k,2}$, where the argument of this partial trace is seen as an operator in $\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})$ by restriction of its input and output to $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, which is a subspace of $\mathcal{H}$ for every $k\in[K]$.
Furthermore, the range of $\mathcal{F}^{*}$ is a *-algebra and the following decompositions hold:
$\operatorname{Rg}(\mathcal{F}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}},$ (10)
$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$ (11)
Note that $\mathcal{F}^{*}$ may not be faithful, indeed the states $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$ are not assumed to be full-rank.
We call the shape of $\mathcal{F}$ the $K$-dimensional vector whose coordinates are the dimensions $(d_{k})_{k\in[K]}$ sorted in non-increasing order, which we write $\lambda(\mathcal{F})$.
As an example of a shape vector, consider the qubit channel $\mathcal{F}:\mathcal{L}({\mathbb{C}}^{2})\rightarrow\mathcal{L}({\mathbb{C}}^{2})$ defined as
$\mathcal{F}:x\mapsto\operatorname{\operatorname{tr}}(x)|0\rangle\!\langle 0|=\begin{pmatrix}\operatorname{\operatorname{tr}}(x)&0\\ 0&0\end{pmatrix}$,
where the matrix expression is written for the canonical base of ${\mathbb{C}}^{2}$. Then, in the notations of eq. (7), we have $\mathcal{H}={\mathbb{C}}^{2}$, $K=1$, $\mathcal{H}_{k,1}={\mathbb{C}}$ and $\mathcal{H}_{k,2}={\mathbb{C}}^{2}$ so that $\lambda(\mathcal{F})=(1)$. Note also that, in this case, $\mathcal{F}$ is not faithful as $|0\rangle\!\langle 0|$ is not full-rank in $\mathcal{L}({\mathbb{C}}^{2})$.
We remark that we have $\lambda(\mathcal{F})=(1)$ if and only if $\mathcal{F}$ is a replacer channel, i.e. $\mathcal{F}:x\mapsto\operatorname{\operatorname{tr}}(x)\rho$ with $\rho$ a fixed state. This fact follows easily from eq. (9) taking $K=1$.
[…]
Proposition 2 (Theorem 2.6 and Algorithm 1 in [7]).
[…]
Proposition 3.
[…]
$\mathcal{F}=\mathcal{D}\circ\widetilde{\mathcal{F}}\circ\mathcal{E}$. (12)
Proof.
Let $\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$ and $\widetilde{\mathcal{H}}=\bigoplus_{k=1}^{K}\widetilde{\mathcal{H}}_{k,1}\otimes\widetilde{\mathcal{H}}_{k,2}$ be such that
[…]
Then, let
$\mathcal{E}:x\in\mathcal{L}(\mathcal{H})\mapsto\sum_{k=1}^{K}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\tilde{\omega}_{k}\text{ and }\mathcal{D}:x\in\mathcal{L}(\widetilde{\mathcal{H}})\mapsto\sum_{k=1}^{K}U_{k}^{*}\widetilde{\operatorname{\operatorname{tr}}}_{k,2}(\widetilde{P}_{k}x\widetilde{P}_{k})U_{k}\otimes\rho_{k}$
with a set of states $\{\widetilde{\omega}_{k}\in\mathcal{D}(\widetilde{\mathcal{H}}_{k,2})~|~k\in[K]\}$ and, for all $k\in[K]$, a unitary $U_{k}:\mathcal{H}_{k,1}\mapsto\widetilde{\mathcal{H}}_{k,1}$. These unitaries allow to identify $\mathcal{L}(\mathcal{H}_{k,1})$ with $\mathcal{L}(\widetilde{\mathcal{H}}_{k,1})$ for each $k\in[K]$. Then, for $x\in\mathcal{L}(\mathcal{H})$, we have for any $k,k^{\prime}\in[K]$
$\widetilde{\operatorname{\operatorname{tr}}}_{k^{\prime},2}(\widetilde{P}_{k^{\prime}}(U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\tilde{\omega}_{k})\widetilde{P}_{k^{\prime}})=\delta_{k=k^{\prime}}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}.$
As a result,
$\widetilde{\mathcal{F}}(\mathcal{E}(x))=\sum_{k=1}^{K}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\widetilde{\rho}_{k}.$
It follows that
$\mathcal{D}(\widetilde{\mathcal{F}}(\mathcal{E}(x)))$ $=\sum_{k=1}^{K}U_{k}^{*}\widetilde{\operatorname{\operatorname{tr}}}_{k,2}(U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\widetilde{\rho}_{k})U_{k}\otimes\rho_{k}$
$=\sum_{k=1}^{K}U_{k}^{*}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}U_{k}\otimes\rho_{k}$
$=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}=\mathcal{F}(x).$
We now show that the range and the shape vector of idempotent channels preserve the tensor product, which property will be used in the proofs of our Theorems 1 and 2.
Proposition 4.
For $\mathcal{F}$ an idempotent channel, the range of $\mathcal{F}^{*}$ and thus the shape of $\mathcal{F}$ satisfy for all $n\in\mathbb{N}^{*}$:
$\operatorname{Rg}(\mathcal{F}^{*\otimes n})=\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n},$ (13)
$\lambda(\mathcal{F}^{\otimes n})=\lambda(\mathcal{F})^{\otimes n}.$ (14)
The rest of this section is devoted to proving Propositions 1 and 4.
We begin by proving Proposition 1. We remark that the set of fixed points and the range of an idempotent function coincide.
Lemma 3.
Let $f:\Omega\rightarrow\Omega$ be an arbitrary idempotent function, with $\Omega$ an arbitrary set, then
$\operatorname{Rg}(f)\coloneqq\{f(x)~|~x\in\Omega\}=\{x\in\Omega~|~f(x)=x\}.$ (15)
Proof.
If $f(x)=x$, then $f(f(x))=f(x)=x$, so that $x\in\operatorname{Rg}(f)$. For the other inclusion, if $y=f(x)$, we also have $f(y)=f(f(x))=f(x)=y$, so $y$ is a fixed point. ∎
Proposition 1 basically follows from statements in [18] and we detail it here only for the convenience of the reader.
Proof of Proposition 1.
[…]
The structure of these algebras is well-understood (see for example [16, 5, 6]). In finite dimensions, any *-algebra can be decomposed as a direct sum of algebras of complex square matrices (which, in the theory of operator algebras, are called factors [16, 5, 6]). This allows us to write
$\operatorname{Rg}(\mathcal{F}^{*})\cong\bigoplus_{k=1}^{K}\mathcal{L}({\mathbb{C}}^{d_{k}}),$ (16)
for a set of integers $\{d_{1},\dots,d_{K}\}$, where $\cong$ denotes a *-algebra isomorphism. Furthermore, this decomposition is unique up to permutation of the terms in the right-hand side of eq. (16) and characterizes uniquely each *-algebra, in the sense that two *-algebras are isomorphic if and only if they have the same decomposition up to permutation of the terms in the direct sum.
We can go even further and give the general form of $\mathcal{F}^{*}$ using Proposition 1.5 in [18]. Indeed, we can decompose the Hilbert space $\mathcal{H}$ as a direct sum $\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, where we write $d_{k}=\dim(\mathcal{H}_{k,1})$ and $m_{k}=\dim(\mathcal{H}_{k,2})$, define a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$ and a set of orthogonal projectors $\{P_{k}~|~\operatorname{Rg}(P_{k})=\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},k\in[K]\}$ such that $\mathcal{F}^{*}$ takes the form:
$\forall x\in\mathcal{L}(\mathcal{H}),~\mathcal{F}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}}.$ (17)
The range $\operatorname{Rg}(\mathcal{F}^{*})$ will then be decomposed as
$\operatorname{Rg}(\mathcal{F}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}.$ (18)
Indeed, let $x=\sum_{k=1}^{K}x_{k}\otimes\mathds{1}_{m_{k}}$ with $x_{k}\in\mathcal{L}(\mathcal{H}_{k,1})$ for all $k\in[K]$, then:
$\mathcal{F}^{*}(x)$ $=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(x_{k}\otimes\rho_{k})\otimes\mathds{1}_{m_{k}}=x,$
so that $x\in\operatorname{Rg}(\mathcal{F}^{*})$. This proves eq. (10). Note that eq. (18) implies eq. (16) as each term $\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}$ is isomorphic, as a *-algebra, to $\mathcal{L}(\mathcal{H}_{k,1})$.
Then, as $\mathcal{F}$ is the adjoint of $\mathcal{F}^{*}$, we can take the adjoint of eq. (17) and get:
$\forall x\in\mathcal{L}(\mathcal{H}),~\mathcal{F}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}.$ (19)
Thus, whenever $\mathcal{F}$ is an idempotent channel, its range admits the decomposition
$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$ (20)
which proves eq. (11). ∎
We then proceed to the proof of Proposition 4, which will allow us to reduce the study of the problem of emulating $k$ copies of $\mathcal{F}$ using $n$ copies of $\mathcal{G}$ for general pairs of integers $(k,n)$ to the case where $k=n$. The fact that we can do such a reduction is probably the reason why we get a single-letter formula for the emulation capacity.
Proof of Proposition 4.
As, by definition, $\lambda(\mathcal{F})$ is obtained from the decomposition in factors of the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$, it follows that eq. (14) is a consequence of eq. (13).
Let $n\in\mathbb{N}^{*}$, we take the $n$-th tensor power of eq. (17) to get that, for $x\in\mathcal{L}(\mathcal{H})$,
$\mathcal{F}^{*\otimes n}(x)=\sum_{\bar{k}\in[K]^{n}}\operatorname{\operatorname{tr}}_{\bar{k},2}(P_{\bar{k}}xP_{\bar{k}}(\mathds{1}_{d_{\bar{k}}}\otimes\rho_{\bar{k}}))\otimes\mathds{1}_{m_{\bar{k}}},$ (21)
with $\bar{k}=(k_{1},\dots,k_{n})$ ranging on all the $n$-uples in $K]^{n}$, $\operatorname{\operatorname{tr}}_{\bar{k},2}=\operatorname{\operatorname{tr}}_{k_{1},2}\otimes\dots\otimes\operatorname{\operatorname{tr}}_{k_{n},2}$ are the product partial traces, $P_{\bar{k}}=P_{k_{1}}\otimes\dots\otimes P_{k_{n}}$ are the product projections, $\rho_{\bar{k}}=\rho_{k_{1}}\otimes\dots\otimes\rho_{k_{n}}$ are product states, and similarly $\mathds{1}_{d_{\bar{k}}}=\mathds{1}_{d_{k_{1}}}\otimes\dots\otimes\mathds{1}_{d_{k_{n}}}$. Hence, the range of $\mathcal{F}^{*}$ can be decomposed as:
$\operatorname{Rg}(\mathcal{F}^{*\otimes n})=\bigoplus_{\bar{k}\in[K]^{n}}\mathcal{L}(\mathcal{H}_{\bar{k}})\otimes\mathds{1}_{m_{\bar{k}}}=\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n},$ (22)
which completes the proof of Proposition 4. ∎
We end this section by the following remark on the origin of the concept of shape vectors.
Remark 1.
Note that the shape of an idempotent channel $\mathcal{F}$ coincides with the shape of the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$. The notion of shape vector for *-algebras was introduced by Kuperberg in [11]. If $\mathcal{A}$ is a *-algebra, we can decompose
$\mathcal{A}\cong\bigoplus_{k=1}^{K}\mathcal{L}({\mathbb{C}}^{d_{k}}),$
for a set of integers $\{d_{k}~|~k\in[K]\}$ and we write $\lambda(\mathcal{A})$ the shape vector of $\mathcal{A}$, which coordinates are the integers $(d_{k})_{k\in[K]}$ sorted in non-increasing order. [^2] In fact, in all this article and in Kuperberg’s [11], the order with respect to which the coordinates of shape vectors are sorted is actually irrelevant.
4 Emulation of idempotent channel with zero error
Now that we have presented the decomposition properties of idempotent channels, we move on to proving Theorem 1 in this section, which is divided into two subsections. In the first one, we prove that every rate lesser than
$\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}$ (23)
is achievable. In the second one, we prove that there is no rate strictly greater than eq. (23).
Before presenting these proofs, we provide some examples for the expression eq. (23). In particular, we gather in Table 2 the expression of the emulation capacity $C(\mathcal{G}\mapsto\mathcal{F})$ when $\mathcal{F}$ or $\mathcal{G}$ is an identity quantum channel or a completely dephasing channel. In fact, when $\mathcal{F}=\mathrm{Id}_{d}$, then $\lambda(\mathcal{F})=(d)$ and thus $\|\lambda(\mathcal{F})\|_{p}=d$ for all $p\geq 1$. As $\|\lambda(\mathcal{G})\|_{p}$ is non-increasing in $p$, this shows that $C(\mathcal{G}\mapsto\mathcal{F})=\frac{\log(\|\lambda(\mathcal{G})\|_{\infty})}{\log(d)}$. The other cases are similar. Note that, when $\mathcal{G}=\Delta_{d}$, $\log(\|\lambda(\Delta_{d})\|_{\infty})=\log(1)=0$. Therefore, in Table 2, we divided the case $\mathcal{G}=\Delta_{d}$ in two more cases: the one where $\mathcal{F}$ is any idempotent channel such that at least one coordinate of $\lambda(\mathcal{F})$ is greater or equal to $2$, and the one where $\lambda(\mathcal{F})=(1,\dots,1)$, in which case, for all $p\in[1,+\infty[$,
$\frac{\log(\|\lambda(\Delta_{d})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}=\frac{p\log(\|\lambda(\Delta_{d})\|_{p})}{p\log(\|\lambda(\mathcal{F})\|_{p})}=\frac{\log(d)}{\log(\|\lambda(\mathcal{F})\|_{1})}.$
Note that in the first case, i.e. $\lambda(\mathcal{F})$ has a coefficient greater or equal to two, the emulation capacity of $\mathcal{F}$ by the completely dephasing channel is zero, that is that it is impossible to emulate a channel preserving some quantum information with a classical one, without error.
$\mathcal{F}$ $\mathcal{G}$ $C(\mathcal{G}\mapsto\mathcal{F})$
$\mathrm{Id}_{d}$ $\mathcal{G}$ $\frac{\log(\|\lambda(\mathcal{G})\|_{\infty})}{\log(d)}$
$\mathcal{F}$ $\mathrm{Id}_{d}$ $\frac{\log(d)}{\log(\|\lambda(\mathcal{F})\|_{1})}$
$\Delta_{d}$ $\mathcal{G}$ $\frac{\log(\|\lambda(\mathcal{G})\|_{1})}{\log(d)}$
 $\mathcal{F}\text{ s.t. }\lambda(\mathcal{F})\neq(1,\dots,1)$ $\Delta_{d}$ 0
 $\mathcal{F}\text{ s.t. }\lambda(\mathcal{F})=(1,\dots,1)$ $\Delta_{d}$ $\frac{\log(d)}{\log(\|\lambda(\mathcal{F})\|_{1})}$
Table 2: Emulation capacities when either $\mathcal{F}$ or $\mathcal{G}$ is the identity or the completely dephasing channel.
Then, we also note that, as a corollary to our Theorem 1, the emulation capacity of idempotent channels is additive under tensor products of source channels, which yields that the zero-error quantum and classical capacities of idempotent channels are additive.
Corollary 1.
Let $\mathcal{F}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be three idempotent channels, then
$C(\mathcal{G}_{1}\otimes\mathcal{G}_{2}\mapsto\mathcal{F})=C(\mathcal{G}_{1}\mapsto\mathcal{F})+C(\mathcal{G}_{2}\mapsto\mathcal{F}).$ (24)
Proof.
By Theorem 1, we have
$C(\mathcal{G}_{1}\otimes\mathcal{G}_{2}\mapsto\mathcal{F})$ $=\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G}_{1}\otimes\mathcal{G}_{2})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}$
$=\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G}_{1})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}+\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G}_{2})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}$
$=C(\mathcal{G}_{1}\mapsto\mathcal{F})+C(\mathcal{G}_{2}\mapsto\mathcal{F}),$
where we used Proposition 4 and the multiplicativity of the $\ell_{p}$-norms over the tensor product in the second equation. ∎
The following example shows that the emulation capacity is not reversible, i.e. that $C(\mathcal{G}\mapsto\mathcal{F})\neq C(\mathcal{F}\mapsto\mathcal{G})^{-1}$ in general.
Example 1.
Let $\mathcal{F}=\mathrm{Id}_{4}$ be the identity channel on $\mathcal{L}({\mathbb{C}}^{4})$ and $\mathcal{G}:\mathcal{L}({\mathbb{C}}^{4})\rightarrow\mathcal{L}({\mathbb{C}}^{4})$ be such that $\lambda(\mathcal{G})=(2,2)$. Then, by the first row of Table 2, we have
$C(\mathcal{G}\mapsto\mathcal{F})=\frac{\log(\|\lambda(\mathcal{G})\|_{\infty})}{\log(4)}=\frac{\log(2)}{\log(4)}=\frac{1}{2},$
whereas by the second row,
$C(\mathcal{F}\mapsto\mathcal{G})=\frac{\log(4)}{\log(\|\lambda(\mathcal{G})\|_{1})}=\frac{\log(4)}{\log(4)}=1.$
Therefore,
$C(\mathcal{G}\mapsto\mathcal{F})^{-1}=2\neq 1=C(\mathcal{F}\mapsto\mathcal{G}).$
This example shows that neither the emulation capacity nor the strong converse rate given in Theorem 2 are reversible, since, in this example, the emulation capacity matches with the strong converse rate of Theorem 2, as $\mathcal{F}$ is taken to be an identity channel.
Although, when either $\mathcal{F}$ or $\mathcal{G}$ is an identity channel or a completely dephasing one, the infimum in eq. (1) is reached for $p\in\{1,+\infty\}$, we show in the following example, that this is not the case in general.
In particular, the fact that the infimum in eq. (1) can be reached for $p\not\in\{1,+\infty\}$ implies that, for general idempotent channels $\mathcal{F}$ and $\mathcal{G}$, the chain of emulation $\mathcal{G}\mapsto\mathrm{Id}_{d}\mapsto\mathcal{F}$ (where each arrow represents an emulation) is less efficient than directly emulating $\mathcal{F}$ by $\mathcal{G}$.
Example 2.
Let $\mathcal{F}:\mathcal{L}({\mathbb{C}}^{8})\rightarrow\mathcal{L}({\mathbb{C}}^{8})$ and $\mathcal{G}:\mathcal{L}({\mathbb{C}}^{15})\rightarrow\mathcal{L}({\mathbb{C}}^{15})$ be two idempotent channels such that $\lambda(\mathcal{F})=(5,3)$ and $\lambda(\mathcal{G})=(10,3,1,1)$, then, we numerically find that the infimum
$C(\mathcal{G}\mapsto\mathcal{F})=\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}\approx 1.29916,$
is reached for $p\approx 1.15401$. In contrast, the infimum
$C(\mathcal{F}\mapsto\mathcal{G})=\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{F})\|_{p})}{\log(\|\lambda(\mathcal{G})\|_{p})}=\frac{\log(5)}{\log(10)},$
is reached for $p=+\infty$. We can verify that
$C(\mathcal{F}\mapsto\mathcal{G})^{-1}=\frac{\log(10)}{\log(5)}\approx 1.43068\neq C(\mathcal{G}\mapsto\mathcal{F}).$
4.1 Achievability in Theorem 1
To prove the achievability direction of Theorem 1, we first show the following proposition in the case where only one copy of $\mathcal{F}$ is emulated for each copy of $\mathcal{G}$ used, we will then use Proposition 4 on the compatibility of the shape vector with the tensor product to conclude.
Proposition 5.
Let $\mathcal{F}$ and $\mathcal{G}$ be two idempotent channels such that, for all $p\in[1,+\infty]$,
$\|\lambda(\mathcal{F})\|_{p}<\|\lambda(\mathcal{G})\|_{p}.$
Then, there exists $n\in\mathbb{N}^{*}$, $\mathcal{E}$ and $\mathcal{D}$ two channels such that
$\mathcal{F}^{\otimes n}=\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}.$
To prove this proposition, we use the following theorem on embeddability of *-algebras.
Theorem 4 (Proven in [11]).
If $\mathcal{A}$ and $\mathcal{B}$ are two finite-dimensional *-algebras such that, for all $p\in[1,+\infty]$,
$\|\lambda(\mathcal{A})\|_{p}<\|\lambda(\mathcal{B})\|_{p},$ (25)
then, there is $n\in\mathbb{N}^{*}$ such that there exists a subunital injective *-homomorphism $\iota:\mathcal{A}^{\otimes n}\rightarrow\mathcal{B}^{\otimes n}$.
Proof.
This theorem, despite not stated explicitly in [11], is proven in the proof of Theorem 1.1 of [11]. ∎
For our achievability proof, it will be useful to lift $\iota$ into a unital completely positive map.
Lemma 5.
Let $\mathcal{A}$, $\mathcal{B}$ be two finite-dimensional *-algebras and $\iota:\mathcal{A}\rightarrow\mathcal{B}$ be a subunital *-homomorphism. Let $\mathds{1}_{\mathcal{A}}\in\mathcal{A}$ and $\mathds{1}_{\mathcal{B}}\in\mathcal{B}$ be the identity operator of $\mathcal{A}$ and $\mathcal{B}$ respectively. Then, for all $x\in\mathcal{A}$,
$\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}})).$ (26)
Furthermore, $P_{\iota}\coloneqq\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector, and the map
$\widetilde{\iota}:x\mapsto\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})$ (27)
is unital completely positive. Finally, under the assumption that $\iota$ is injective, we write $\iota^{-1}:\iota(\mathcal{A})\subseteq\mathcal{B}\rightarrow\mathcal{A}$ its partial inverse, then
$\widetilde{\iota}^{-1}:x\in\mathcal{B}\mapsto\iota^{-1}(P_{\iota}xP_{\iota})$ (28)
is a unital completely positive map such that
$\forall x\in\mathcal{A},~\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)=x.$ (29)
Note that none of the two maps $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are *-homomorphisms unless $\iota$ is unital.
Proof.
We start by proving eq. (26). Let $x\in\mathcal{A}$, we can use the polarisation identity (eq. 1.45 in [18]) to write
$x=\frac{1}{4}\sum_{k=0}^{3}i^{k}(x+i^{k}\mathds{1}_{\mathcal{A}})^{*}(x+i^{k}\mathds{1}_{\mathcal{A}}),$
so that $x$ is decomposed as a linear combination of four positive operators in $\mathcal{A}$.
Thus, it suffices to show eq. (26) for positive semidefinite operators. For positive semidefinite $x\in\mathcal{A}$,
$0\leq x\leq\|x\|\mathds{1}_{\mathcal{A}}.$
Since $\iota$ is completely positive,
$0\leq\iota(x)\leq\|x\|\iota(\mathds{1}_{\mathcal{A}}).$
Thus, $\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$. Now,
$\iota(\mathds{1}_{\mathcal{A}})^{2}=\iota(\mathds{1}_{\mathcal{A}}^{2})=\iota(\mathds{1}_{\mathcal{A}}),$
and, as $\iota$ is a *-homomorphism, it preserves the *-operation on $\mathcal{A}$, so that
$\iota(\mathds{1}_{\mathcal{A}})^{*}=\iota(\mathds{1}_{\mathcal{A}}).$
This shows that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector in $\mathcal{B}$.
For $\widetilde{\iota}$, note that, as $\iota$ is subunital, $P_{\iota}\leq\mathds{1}_{\mathcal{B}}$, so that $\mathds{1}_{\mathcal{B}}-P_{\iota}\geq 0$ and $x\mapsto\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})$ is completely positive. The unitality is trivial.
Then, as $\iota^{-1}$ is a *-homomorphism, it is completely positive. Furthermore, as $P_{\iota}$ is an orthogonal projector, $x\in\mathcal{B}\mapsto P_{\iota}xP_{\iota}$ is completely positive. Therefore the composition $x\mapsto\iota^{-1}(P_{\iota}xP_{\iota})$ is completely positive as well. Then $\widetilde{\iota}^{-1}(\mathds{1}_{\mathcal{B}})=\iota^{-1}(P_{\iota})=\iota^{-1}(\iota(\mathds{1}_{\mathcal{A}}))=\mathds{1}_{\mathcal{A}}$.
Finally, for $x\in\mathcal{A}$, we have
$\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)$ $=\iota^{-1}(P_{\iota}\bigl(\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})\bigr)P_{\iota})$
$=\iota^{-1}(P_{\iota}\iota(x)P_{\iota})$
$=x,$
where we used for the last equality that $\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(P_{\iota})$ so that $P_{\iota}\iota(x)P_{\iota}=\iota(x)$. ∎
We now prove Proposition 5.
Proof of Proposition 5.
[…]
$\|\lambda(\operatorname{Rg}(\mathcal{F}^{*}))\|_{p}=\|\lambda(\mathcal{F})\|_{p}<\|\lambda(\mathcal{G})\|_{p}=\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}.$
Therefore, by Theorem 4, there is an $n\in\mathbb{N}^{*}$ such that there exists an injective subunital *-homomorphism $\iota:\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n}\rightarrow\operatorname{Rg}(\mathcal{G}^{*})^{\otimes n}$ between the two *-algebras $\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n}$ and $\operatorname{Rg}(\mathcal{G}^{*})^{\otimes n}$. Then, by Proposition 4, $\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n}=\operatorname{Rg}(\mathcal{F}^{*\otimes n})$ and $\operatorname{Rg}(\mathcal{G}^{*})^{\otimes n}=\operatorname{Rg}(\mathcal{G}^{*\otimes n})$, so that $\iota$ injects $\operatorname{Rg}(\mathcal{F}^{*\otimes n})$ into $\operatorname{Rg}(\mathcal{G}^{*\otimes n})$. We then show that we have the factorisation:
$\mathcal{F}^{*\otimes n}=\widetilde{\iota}^{-1}\circ\mathcal{G}^{*\otimes n}\circ\widetilde{\iota}\circ\mathcal{F}^{*\otimes n},$ (30)
with $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ being defined as in Lemma 5. Indeed, for $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$,
$\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\iota(\mathcal{F}^{*\otimes n}(x))+\frac{\operatorname{\operatorname{tr}}(\mathcal{F}^{*\otimes n}(x))}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{F}^{\otimes n}})}(\mathds{1}_{\mathcal{G}^{\otimes n}}-\iota(\mathds{1}_{\mathcal{F}^{\otimes n}}))\in\operatorname{Rg}(\mathcal{G}^{*\otimes n}),$
where $\mathds{1}_{\mathcal{F}^{\otimes n}}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$ and $\mathds{1}_{\mathcal{G}^{\otimes n}}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$. Since $\mathcal{G}^{*\otimes n}$ is idempotent, every element of its range is one of its fixed points, so that $\mathcal{G}^{*\otimes n}(\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x)))=\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))$. We finally have by eq. (29) that $\widetilde{\iota}^{-1}\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\mathcal{F}^{*\otimes n}(x)$.
We then write $\mathcal{E}^{*}=\widetilde{\iota}^{-1}$ and $\mathcal{D}^{*}=\widetilde{\iota}\circ\mathcal{F}^{*\otimes n}$. By Lemma 5, both $\mathcal{E}^{*}$ and $\mathcal{D}^{*}$ are unital completely positive, so that their adjoint $\mathcal{E}$ and $\mathcal{D}$ are channels. Taking the adjoint of eq. (30), we finally get
$\mathcal{F}^{\otimes n}=\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}.$
Then we can prove the achievability in Theorem 1.
Proof of achievability in Theorem 1.
Let $\mathcal{F}$, $\mathcal{G}$ be idempotent channels and
$r<\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}.$
By density of the rational numbers into the reals, there exists $k,n\in\mathbb{N}^{*}$ such that
$r\leq\frac{k}{n}<\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}.$
Then, for every $p\in[1,+\infty]$, we have
$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}<\|\lambda(\mathcal{G}^{\otimes n})\|_{p},$
where we used Proposition 4 on the shape vectors of the tensor product of channels.
Then, by Proposition 5, there is $m\in\mathbb{N}^{*}$, $\mathcal{E}$ and $\mathcal{D}$ channels such that
$\mathcal{F}^{\otimes km}=\mathcal{D}\mathcal{G}^{\otimes nm}\mathcal{E},$
and $r\leq k/n=km/(nm)$ is thus an achievable rate. ∎
4.2 Converse of Theorem 1
We now move on to the proof of the converse of Theorem 1. To prove this direction, we use a result of [11] which is almost a converse to Theorem 4.
Theorem 6 (Forward direction of Theorem 1.1 in [11]).
Let $\mathcal{A}$ and $\mathcal{B}$ be two finite-dimensional *-algebras. If there is an injective *-homomorphism of $\mathcal{A}$ into $\mathcal{B}$, then for all $p\in[1,+\infty]$,
$\|\lambda(\mathcal{A})\|_{p}\leq\|\lambda(\mathcal{B})\|_{p}.$ (31)
Note that Theorems 4 and 6 are not converse one from another because the inequality between the $\ell_{p}$-norms of the shape vectors is strict in one case but not in the other.
The converse of Theorem 1 then follows from Theorem 6 combined with the following proposition.
Proposition 6.
[…]
$\mathcal{F}=\mathcal{D}\mathcal{G}\mathcal{E}.$ (32)
We write $\mathds{1}_{\mathcal{F}}$ and $\mathds{1}_{\mathcal{G}}$ the identity operator in $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$, respectively. To simplify our notations, we also denote $e_{\mathcal{G}\mathcal{E}}=\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})^{0}$. Define
$\widetilde{\mathcal{G}}^{*}:x\mapsto e_{\mathcal{G}\mathcal{E}}\mathcal{G}^{*}(x)e_{\mathcal{G}\mathcal{E}}.$ (33)
Then, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism which embeds $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ as a subalgebra.
Furthermore, for all $p\in[1,+\infty]$
$\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}.$
The proof of this proposition uses the following lemma.
Lemma 7.
Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be a quantum channel and write $e=\Phi(\mathds{1})^{0}$. Then
  1. i)
    For all $x\in\mathcal{L}(\mathcal{H}^{\prime})$, $\Phi^{*}(x)=\Phi^{*}(exe)$.
  2. ii)
    For all positive semidefinite $x\in\mathcal{L}(\mathcal{H}^{\prime})$, $\Phi^{*}(x)=0$ implies $exe=0$.
Proof.
Let $y\in\mathcal{L}(\mathcal{H})$ be positive semidefinite. By positivity of $\Phi$,
$\Phi(y)\leq\|y\|\Phi(\mathds{1}).$
Therefore, $\operatorname{\operatorname{supp}}(\Phi(y))\subseteq\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))=e$. Then, for all $x\in\mathcal{L}(\mathcal{H}^{\prime})$,
$\operatorname{\operatorname{tr}}(y\Phi^{*}(x))=\operatorname{\operatorname{tr}}(\Phi(y)x)=\operatorname{\operatorname{tr}}(e\Phi(y)ex)=\operatorname{\operatorname{tr}}(y\Phi^{*}(exe)).$
By linearity, this equation extends to all $y\in\mathcal{L}(\mathcal{H})$. Thus, $\Phi^{*}(x)=\Phi^{*}(exe)$.
For the second statement, let $x\in\mathcal{L}(\mathcal{H}^{\prime})$ be positive semidefinite such that $\Phi^{*}(x)=0$. Then
$0=\operatorname{\operatorname{tr}}(\Phi^{*}(x))=\operatorname{\operatorname{tr}}(\Phi(\mathds{1})x)=\operatorname{\operatorname{tr}}(e\Phi(\mathds{1})ex)\geq\lambda_{\min}(\Phi(\mathds{1}))\operatorname{\operatorname{tr}}(exe),$
with $\lambda_{\min}(\Phi(\mathds{1}))$ the minimal non-zero eigenvalue of $\Phi(\mathds{1})$. Therefore, $\operatorname{\operatorname{tr}}(exe)=0$ and, by positivity of $x$, $exe=0$. ∎
[…]
$\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(x)\leq\mathcal{F}^{*}(x^{*}x).$ (34)
[…]
$\{x\in\mathcal{L}(\mathcal{H}^{\prime})~|~\forall y\in\mathcal{L}(\mathcal{H}^{\prime}),~\mathcal{F}^{*}(xy)=\mathcal{F}^{*}(x)\mathcal{F}^{*}(y)\text{ and }\mathcal{F}^{*}(yx)=\mathcal{F}^{*}(y)\mathcal{F}^{*}(x)\}.$
If $\mathcal{F}^{*}$ is furthermore Hermitian preserving then its multiplicative domain is closed under taking the adjoint and is therefore a *-algebra. Moreover, in this case, $\mathcal{F}^{*}$ acts as a (non necessarily unital) *-homomorphism on its multiplicative domain. Furthermore, when $\mathcal{F}^{*}$ satisfies the Kadison-Schwarz inequality, by Theorem 5.7 in [18], we can write its multiplicative domain as
$\{x\in\mathcal{L}(\mathcal{H}^{\prime})~|~\mathcal{F}^{*}(x^{*}x)=\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(x)\text{ and }\mathcal{F}^{*}(xx^{*})=\mathcal{F}^{*}(x)\mathcal{F}^{*}(x^{*})\}.$
In particular, if $\mathcal{F}^{*}$ is subunital completely positive [^3] In fact, it suffices for $\mathcal{F}^{*}$ to be subunital $2$-positive for it to satisfy the Kadison-Schwarz inequality, see Section 5.2 of [18] for more details., it satisfies the Kadison-Schwarz inequality (eq. (34)) and we can use this second characterisation of its multiplicative domain.
Proof of Proposition 6.
[…]
$\mathcal{F}=\mathcal{D}\mathcal{G}\mathcal{E}.$
Taking the adjoint of this equation, we get
$\mathcal{F}^{*}=\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}=\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{G}^{*}\mathcal{D}^{*},$
where the right-hand side follows from the idempotence of $\mathcal{G}^{*}$. Let $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}})$, applying Lemma 7 to the channel $\Phi=\mathcal{G}\mathcal{E}$, we get that
$\mathcal{F}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}(e_{\mathcal{G}\mathcal{E}}\mathcal{G}^{*}\mathcal{D}^{*}(x)e_{\mathcal{G}\mathcal{E}})=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x),$
with $\widetilde{\mathcal{G}}^{*}$ defined in the statement of Proposition 6. Let $x\in\operatorname{Rg}(\mathcal{F}^{*})$. As $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-algebra, $x^{*}x\in\operatorname{Rg}(\mathcal{F}^{*})$, and as $\mathcal{F}^{*}$ is idempotent, by Lemma 3, we have:
$x^{*}x$ $=\mathcal{F}^{*}(x^{*}x)$
$=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)$
$\overset{(a)}{\geq}\mathcal{E}^{*}\mathcal{G}^{*}(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x))$
$\overset{(b)}{\geq}\mathcal{E}^{*}\mathcal{G}^{*}(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}))\mathcal{E}^{*}\mathcal{G}^{*}(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x))$
$=x^{*}x,$
where we used the Kadison-Schwarz inequality to get both $(a)$ and $(b)$. Since the left-hand side and the right-hand side of this chain of inequalities coincide, we conclude that $(a)$ and $(b)$ are in fact equalities. Therefore,
$\mathcal{E}^{*}\mathcal{G}^{*}(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x))=0,$
where, by the Kadison-Schwarz inequality,
$\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\geq 0.$
Therefore, by the second statement in Lemma 7 applied to the channel $\Phi=\mathcal{G}\mathcal{E}$, we obtain
$\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)=0.$
We can swap the roles of $x$ and $x^{*}$ in $(a)$ and $(b)$ to find that
$\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(xx^{*})-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})=0.$
Therefore, $x$ is in the multiplicative domain of $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$. Thus, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism into the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$. It is injective as, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
$x=\mathcal{F}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x).$
It is unital as it maps the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ to $e_{\mathcal{G}\mathcal{E}}$ which is the identity operator of the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$. Indeed, as both $\mathcal{G}$ and $\mathcal{D}$ are channels, $\mathcal{G}^{*}$ and $\mathcal{D}^{*}$ are unital so that $\mathcal{G}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=\mathds{1}_{\mathcal{G}}$ and $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=e_{\mathcal{G}\mathcal{E}}$.
Now, to show that for all $p\in[1,+\infty]$, $\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}$, it suffices to remark that, by Proposition 1, we can decompose the Hilbert space $\mathcal{H}_{\mathcal{G}}$ as
$\mathcal{H}_{\mathcal{G}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$
and the *-algebra $\operatorname{Rg}(\mathcal{G}^{*})$ as
$\operatorname{Rg}(\mathcal{G}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}},$
with $m_{k}=\dim(\mathcal{H}_{k,2})$ for all $k\in[K]$. Then $\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})\in\operatorname{Rg}(\mathcal{G})$ can itself be decomposed as
$\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})=\sum_{k=1}^{K}x_{k}\otimes\rho_{k},$
with $\{x_{k}\in\mathcal{L}(\mathcal{H}_{k,1})~|~k\in[K]\}$ and $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$. Thus,
$e_{\mathcal{G}\mathcal{E}}=\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})^{0}=\sum_{k=1}^{K}x_{k}^{0}\otimes\rho_{k}^{0},$
and
$e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}=\bigoplus_{k=1}^{K}x_{k}^{0}\mathcal{L}(\mathcal{H}_{k,1})x_{k}^{0}\otimes\rho_{k}^{0}.$
Hence, for all $k\in[K]$, the $k$-th coordinate of the shape vector of $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ satisfies
$\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})_{k}\leq\lambda(\operatorname{Rg}(\mathcal{G}^{*}))_{k},$
which allows us to conclude. ∎
We can now prove the converse of Theorem 1.
Converse of Theorem 1.
[…]
$\mathcal{F}^{\otimes k}=\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}.$
Then by Proposition 6, there is an injective unital *-homomorphism from $\operatorname{Rg}(\mathcal{F}^{*\otimes k})$ into the *-algebra $e_{\mathcal{G}^{\otimes n}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*\otimes n})e_{\mathcal{G}^{\otimes n}\mathcal{E}}$. Then, by Theorem 6 and the second statement of Proposition 6, we have for all $p\in[1,+\infty]$:
$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}=\|\lambda(\operatorname{Rg}(\mathcal{F}^{*\otimes k}))\|_{p}\leq\|\lambda(e_{\mathcal{G}^{\otimes n}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*\otimes n})e_{\mathcal{G}^{\otimes n}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*\otimes n}))\|_{p}.$
By the multiplicativity property of the shape vectors and the $\ell_{p}$-norms with the tensor product (Proposition 4), we get for all $p\in[1,+\infty]$:
$\|\lambda(\mathcal{F})\|_{p}^{k}\leq\|\lambda(\mathcal{G})\|^{n}_{p}.$
Thus,
$\frac{k}{n}\leq\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})},$
which ends the proof. ∎
Finally, we give a slightly stronger version of the converse of Theorem 1 when allowing for dimension dependent errors in the Appendix. This stronger result is an application of approximate $C^{*}$-algebras, which were recently introduced by Kitaev in [9].
5 Channel emulation with errors
In Theorem 1, the converse states that it is impossible to emulate an idempotent channel $\mathcal{F}$ by another idempotent channel $\mathcal{G}$ at a rate higher than the infimum in eq. (1) without error. In this section, we prove Theorem 2 which shows that we can not achieve better rates than
$\inf_{p\in\{1,+\infty\}}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})},$
even if one allows errors. This strong converse rate is probably not tight in general, as the emulation capacity is expressed in Theorem 1 as an infimum over all $p\in[1,+\infty]$, we therefore conjecture that our strong converse rate can be improved to the infimum over all $p\in[1,+\infty]$.
Theorem 2 is actually a corollary of the following theorem, which gives a lower bound on the minimal error achievable for the one-shot emulation of $\mathcal{F}$ by $\mathcal{G}$. Note that this theorem can be seen as a generalisation of eq. (18) in [10] about emulation of identity channels. In fact, even in the special case of identity channels, we strengthen the bound of [10] by a factor of $2$.
Theorem 8.
Let $\mathcal{F}$ and $\mathcal{G}$ be two idempotent channels, then
$\frac{1}{2}\inf_{\mathcal{D},\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\geq 1-\min_{p\in\{1,+\infty\}}\frac{\|\lambda(\mathcal{G})\|_{p}}{\|\lambda(\mathcal{F})\|_{p}}.$ (35)
To prove this theorem, we use the Holevo-Helstrom bound applied to the problem of channel discrimination.
Lemma 9 (see e.g. Theorem 3.52 in [17]).
Let $\Phi_{1},~\Phi_{2}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be two quantum channels. For any choice of auxiliary system $\mathcal{L}(\mathcal{H}^{\prime\prime})$ of dimension $d=\dim(\mathcal{H}^{\prime\prime})$, for any choice of positive semidefinite operator $\mu\in\mathcal{L}(\mathcal{H}^{\prime}\otimes\mathcal{H}^{\prime\prime})$ satisfying $\mu\leq\mathds{1}$ and of density operator $\sigma\in\mathcal{D}(\mathcal{H}\otimes\mathcal{H}^{\prime\prime})$,
$\operatorname{\operatorname{tr}}(\mu\Phi_{1}\otimes\mathrm{Id}_{d}(\sigma))-\operatorname{\operatorname{tr}}(\mu\Phi_{2}\otimes\mathrm{Id}_{d}(\sigma))\leq\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}.$ (36)
We can now prove Theorem 8.
Proof of Theorem 8.
[…]
$\frac{1}{2}\inf_{\mathcal{D},\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\geq 1-\frac{\|\lambda(\mathcal{G})\|_{1}}{\|\lambda(\mathcal{F})\|_{1}},$ (37)
and
$\frac{1}{2}\inf_{\mathcal{D},\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\geq 1-\frac{\|\lambda(\mathcal{G})\|_{\infty}}{\|\lambda(\mathcal{F})\|_{\infty}},$ (38)
which taken together amount to the statement of Theorem 8.
To prove each equation, we use Lemma 9 on the channels $\Phi_{1}=\mathcal{F}$ and $\Phi_{2}=\mathcal{D}\mathcal{G}\mathcal{E}$ and we find an Ansatz $(\sigma,\mu)$ in eq. (36) such that on the one hand
$\operatorname{\operatorname{tr}}(\mu\mathcal{F}\otimes\mathrm{Id}_{d}(\sigma))=1,$
and, on the other hand
$\operatorname{\operatorname{tr}}(\mu\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}_{d}(\sigma))\leq\frac{\|\lambda(\mathcal{G})\|_{p}}{\|\lambda(\mathcal{F})\|_{p}},$
with $p=1$ for one Ansatz and $p=+\infty$ for the other.
We begin with the case $p=1$. As we saw in Proposition 1, we can decompose $\mathcal{H}_{\mathcal{F}}$ as a direct sum
$\mathcal{H}_{\mathcal{F}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$
so that
$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
[…]
$\sigma=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}\otimes|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}},$ (39)
where $|\nu\rangle\!\langle\nu|_{k}=|\nu\rangle_{k}\!\langle\nu|_{k}$ for notational simplicity. Then, we take
$\mu=\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\otimes|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}.$ (40)
We now have:
$\operatorname{\operatorname{tr}}(\mu\mathcal{F}\otimes\mathrm{Id}(\sigma))$
$=\operatorname{\operatorname{tr}}(\mathcal{F}^{*}\otimes\mathrm{Id}(\mu)\sigma)$
$=\operatorname{\operatorname{tr}}(\mu\sigma)$
$=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\operatorname{\operatorname{tr}}(\bigl(\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\otimes|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\bigr)\bigl(\sum_{k^{\prime}=1}^{K}\sum_{\nu^{\prime}=1}^{d_{k^{\prime}}}|\nu^{\prime}\rangle\!\langle\nu^{\prime}|_{k^{\prime}}\otimes\frac{\mathds{1}_{m_{k^{\prime}}}}{m_{k^{\prime}}}\otimes|\nu^{\prime}\rangle\!\langle\nu^{\prime}|_{k^{\prime}}\otimes\frac{\mathds{1}_{m_{k^{\prime}}}}{m_{k^{\prime}}}\bigr))$
$=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\operatorname{\operatorname{tr}}(\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}|\nu\rangle\!\langle\nu|_{k}\otimes|\nu\rangle\!\langle\nu|_{k})$
$=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\sum_{k=1}^{K}d_{k}$
$=1,$
where, for the second equality, we used the fact that as $|\nu\rangle\!\langle\nu|_{k}\in\mathcal{L}(\mathcal{H}_{k,1})$, $|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\in\operatorname{Rg}(\mathcal{F}^{*})$ for all $k\in[K]$ so that $\mathcal{F}^{*}\otimes\mathrm{Id}(\mu)=\mu$. For the upper bound on $\operatorname{\operatorname{tr}}(\mu\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))$, we have
$\operatorname{\operatorname{tr}}(\mu\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))$
$=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\operatorname{\operatorname{tr}}(\Bigl(\sum_{\begin{aligned} k&\in[K]\\ \nu&\in[d_{k}]\end{aligned}}|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\otimes|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\Bigr)\Bigl(\sum_{\begin{aligned} k^{\prime}&\in[K]\\ \nu^{\prime}&\in[d_{k^{\prime}}]\end{aligned}}\mathcal{D}\mathcal{G}\mathcal{E}\bigl(|\nu^{\prime}\rangle\!\langle\nu^{\prime}|_{k^{\prime}}\otimes\frac{\mathds{1}_{m_{k^{\prime}}}}{m_{k^{\prime}}}\bigr)\otimes|\nu^{\prime}\rangle\!\langle\nu^{\prime}|_{k^{\prime}}\otimes\frac{\mathds{1}_{m_{k^{\prime}}}}{m_{k^{\prime}}}\Bigr)),$
so that
$\operatorname{\operatorname{tr}}(\mu\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\operatorname{\operatorname{tr}}(\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})(\mathcal{D}\mathcal{G}\mathcal{E}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))).$ (41)
From Propososition 1, we can decompose $\mathcal{H}_{\mathcal{G}}$ and $\mathcal{G}$ as
$\mathcal{H}_{\mathcal{G}}=\bigoplus_{k=1}^{K^{\prime}}\mathcal{H}^{\prime}_{k,1}\otimes\mathcal{H}^{\prime}_{k,2},$
and
$\mathcal{G}:x\mapsto\sum_{k=1}^{K^{\prime}}\operatorname{\operatorname{tr}}_{k,2}(P^{\prime}_{k}xP^{\prime}_{k})\otimes\rho^{\prime}_{k},$
with $P^{\prime}_{k}$ the orthogonal projector on $\mathcal{H}^{\prime}_{k,1}\otimes\mathcal{H}^{\prime}_{k,2}$, $\operatorname{\operatorname{tr}}_{k,2}(\cdot)$ the partial trace over $\mathcal{H}^{\prime}_{k,2}$ and $\rho^{\prime}_{k}\in\mathcal{D}(\mathcal{H}^{\prime}_{k,2})$ for all $k\in[K^{\prime}]$. We can therefore factorise $\mathcal{G}$ into two channels, one being the sum of the partial traces and the other being the sum of tensor products with the states $\rho^{\prime}_{k}$ for $k\in[K^{\prime}]$. We write $\mathcal{H}_{\mathcal{G}_{1}}=\bigoplus_{k=1}^{K^{\prime}}\mathcal{H}^{\prime}_{k,1}$ and we define $\mathcal{G}_{1}:\mathcal{L}(\mathcal{H}_{\mathcal{G}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})$ and $\mathcal{G}_{2}:\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ as
$\mathcal{G}_{1}:x$ $\mapsto\sum_{k=1}^{K^{\prime}}\operatorname{\operatorname{tr}}_{k,2}(P^{\prime}_{k}xP^{\prime}_{k}),$ (42)
$\mathcal{G}_{2}:x$ $\mapsto\sum_{k=1}^{K^{\prime}}\widetilde{P}^{\prime}_{k}x\widetilde{P}^{\prime}_{k}\otimes\rho^{\prime}_{k},$ (43)
where $\widetilde{P}^{\prime}_{k}\in\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})$ is the orthogonal projector on $\mathcal{H}^{\prime}_{k,1}$ for all $k\in[K^{\prime}]$. Now, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are channels and we can write $\mathcal{G}=\mathcal{G}_{2}\circ\mathcal{G}_{1}$. Furthermore, let $\mathds{1}_{\mathcal{G}_{1}}$ be the identity operator in $\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})$, we then have, similarly as in the proof of eq. (18) in [10],
$\|\lambda(\mathcal{G})\|_{1}$ $=\sum_{k=1}^{K^{\prime}}\dim(\mathcal{H}^{\prime}_{k,1})$
$=\dim(\mathcal{H}_{\mathcal{G}_{1}})$
$=\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{G}_{1}})$
$\overset{(a)}{=}\operatorname{\operatorname{tr}}(\mathcal{G}_{2}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}}))$
$=\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}\operatorname{\operatorname{tr}}(\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}))$
$\overset{(b)}{\geq}\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}\operatorname{\operatorname{tr}}(\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})\mathcal{G}_{1}\mathcal{E}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))$
$=\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}\operatorname{\operatorname{tr}}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}}\mathcal{D}\mathcal{G}_{2}\mathcal{G}_{1}\mathcal{E}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))$
$=\operatorname{\operatorname{tr}}(\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})(\mathcal{D}\mathcal{G}\mathcal{E}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))),$
where $(a)$ follows from the fact that $\mathcal{D}\mathcal{G}_{2}:\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ is a channel, so that $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is unital. Then $(b)$ follows from the fact that for all $k\in[K]$ and $\nu\in[d_{k}]$, $|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}$ is a state so that $\operatorname{\operatorname{tr}}(\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))=1$, hence $\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}})\leq\mathds{1}_{\mathcal{G}_{1}}$, and that as $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is completely positive, $\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})\geq 0$. We then simply use that if $A,B\geq 0$ with $B\leq\mathds{1}$, $\operatorname{\operatorname{tr}}(A)\geq\operatorname{\operatorname{tr}}(AB)$. When injecting the last equation into eq. (41), we find
$\operatorname{\operatorname{tr}}(\mu\mathcal{D}\mathcal{G}\mathcal{E}(\sigma))\leq\frac{\|\lambda(\mathcal{G})\|_{1}}{\|\lambda(\mathcal{F})\|_{1}}.$
We now move on to the proof of eq. (38), i.e. the case $p=+\infty$. To deal with the case where the blocks of the range of $\mathcal{F}^{*}$ would have multiplicities, we factorise $\mathcal{F}=\mathcal{F}_{2}\circ\mathcal{F}_{1}$ as we did for $\mathcal{G}$ in eq. (42), so that we have the decompositions
$\mathcal{H}$ $=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$
$\mathcal{F}:x$ $\mapsto\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k},$
$\mathcal{F}_{1}:x$ $\mapsto\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}),$
$\mathcal{F}_{2}:x$ $\mapsto\sum_{k=1}^{K}\widetilde{P}_{k}x\widetilde{P}_{k}\otimes\rho_{k},$
where, for $k\in[K]$, $\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})$, $P_{k}$ is the orthogonal projector on $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$ and $\widetilde{P}_{k}\in\mathcal{L}(\mathcal{H}_{\mathcal{F}_{1}})$ is the orthogonal projector on $\mathcal{H}_{k,1}$ with $\mathcal{H}_{\mathcal{F}_{1}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}$. Then, although we have the factorisation $\mathcal{F}=\mathcal{F}_{2}\circ\mathcal{F}_{1}$, we also have the identity $\mathcal{F}_{1}=\mathcal{F}_{1}\circ\mathcal{F}$. Therefore, we can write
$\|\mathcal{F}_{1}-\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}_{1}\circ(\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E})\|_{\diamond}\leq\|\mathcal{F}_{1}\|_{\diamond}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond},$
where the last equality follows from the fact that $\mathcal{F}_{1}$ is a channel and thus, $\|\mathcal{F}_{1}\|_{\diamond}=1$. Therefore, to prove eq. (38), we will prove that
$\|\mathcal{F}_{1}-\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\geq 1-\frac{\|\lambda(\mathcal{G})\|_{\infty}}{\|\lambda(\mathcal{F})\|_{\infty}},$
[…]
$\mu=\sigma=\frac{1}{d}\sum_{\nu,\eta=1}^{d}|\nu\nu\rangle\!\langle\eta\eta|.$ (44)
Then,
$\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\otimes\mathrm{Id}(\sigma))$ $=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\Bigl(\sum_{\nu,\eta=1}^{d}|\nu\nu\rangle\!\langle\eta\eta|\Bigr)\Bigl(\sum_{\nu^{\prime},\mu^{\prime}=1}^{d}\mathcal{F}_{1}(|\nu^{\prime}\rangle\!\langle\eta^{\prime}|)\otimes|\nu^{\prime}\rangle\!\langle\eta^{\prime}|\Bigr))$
$=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\sum_{\nu,\eta,\nu^{\prime},\eta^{\prime}=1}^{d}|\nu\nu\rangle\!\langle\eta\eta||\nu^{\prime}\nu^{\prime}\rangle\!\langle\eta^{\prime}\eta^{\prime}|)$
$=\frac{1}{d^{2}}\bigl(\sum_{\eta,\nu^{\prime}=1}^{d}\langle\eta\eta|\nu^{\prime}\nu^{\prime}\rangle\bigr)\bigl(\sum_{\nu,\eta^{\prime}=1}^{d}\langle\eta^{\prime}\eta^{\prime}|\nu\nu\rangle\bigr)$
$=1,$
where we used the fact that as $|\nu^{\prime}\rangle\!\langle\eta^{\prime}|\in\mathcal{L}(\mathcal{H}_{k_{\infty},1})$ for all $\nu^{\prime},\eta^{\prime}\in[d]$, we have $\mathcal{F}_{1}(|\nu^{\prime}\rangle\!\langle\eta^{\prime}|)=|\nu^{\prime}\rangle\!\langle\eta^{\prime}|$.
For the upper bound on $\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))$, we use techniques much similar to those of the proof of Theorem 2.3 in [7]. First, we remark that the quantity $\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))$ is actually very similar to the entanglement fidelity of $\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}$ and can be easily expressed in terms of its Kraus operators, which we write, for now, $\{K_{i}~|~i\in[R]\}$ [^4] We do not suppose this Kraus decomposition to be minimal.. Then:
$\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))$ $=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\Bigl(\sum_{\nu,\eta=1}^{d}|\nu\nu\rangle\!\langle\eta\eta|\Bigr)\Bigl(\sum_{i=1}^{R}\sum_{\nu^{\prime},\eta^{\prime}=1}^{d}K_{i}|\nu^{\prime}\rangle\!\langle\eta^{\prime}|K_{i}^{*}\otimes|\nu^{\prime}\rangle\!\langle\eta^{\prime}|\Bigr))$
$=\frac{1}{d^{2}}\sum_{i=1}^{R}\sum_{\nu,\eta=1}^{d}\operatorname{\operatorname{tr}}(|\nu\rangle\!\langle\eta|K_{i}|\eta\rangle\!\langle\nu|K_{i}^{*})$
[…]
$=\frac{1}{d^{2}}\sum_{i=1}^{R}|\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}K_{i})|^{2},$
where we write $\widetilde{P}_{\infty}=\widetilde{P}_{k_{\infty}}=\sum_{\nu}|\nu\rangle\!\langle\nu|\in\mathcal{L}(\mathcal{H}_{\mathcal{F}_{1}})$ the orthogonal projector onto $\mathcal{H}_{k_{\infty},1}$. Now, we use the fact that the channel is of the form $\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}$. First, we factorise $\mathcal{G}=\mathcal{G}_{2}\circ\mathcal{G}_{1}$ using the notations of eq. (42) and we furthermore define the pinching map $\mathcal{P}:\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})$ as
$\mathcal{P}:x\mapsto\sum_{k=1}^{K^{\prime}}\widetilde{P}^{\prime}_{k}x\widetilde{P}^{\prime}_{k},$ (45)
with $\widetilde{P}^{\prime}_{k}\in\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})$ the orthogonal projector on $\mathcal{H}^{\prime}_{k,1}$ for all $k\in[K^{\prime}]$. We then have the full factorisation $\mathcal{G}=\mathcal{G}_{2}\circ\mathcal{P}\circ\mathcal{G}_{1}$. For notational simplicity, we write respectively $\{D_{a}\}_{a}$, $\{E_{c}\}_{c}$ the Kraus operators of the channels $\mathcal{F}_{1}\circ\mathcal{D}\circ\mathcal{G}_{2}$ and $\mathcal{G}_{1}\circ\mathcal{E}$. Hence, we can write Kraus operators $\{K_{i}\}_{i}$ of $\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}$ as $\{D_{a}\widetilde{P}_{b}^{\prime}E_{c}\}_{a,b,c}$. Note that we can express $\|\lambda(\mathcal{G})\|_{\infty}$ in terms of the Kraus operators of $\mathcal{P}$ as $\|\lambda(\mathcal{G})\|_{\infty}=\max_{b}\operatorname{\operatorname{tr}}(\widetilde{P}^{\prime}_{b})$. Therefore, we have
$\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))=\frac{1}{d^{2}}\sum_{a,b,c}|\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}D_{a}\widetilde{P}^{\prime}_{b}E_{c})|^{2}.$
By the Cauchy-Schwarz inequality on operators, for $x$ and $y$ of adequate dimensions, we have $|\operatorname{\operatorname{tr}}(x^{*}y)|^{2}\leq\operatorname{\operatorname{tr}}(x^{*}x)\operatorname{\operatorname{tr}}(y^{*}y)$. Hence,
$\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\mathcal{D}\mathcal{G}\mathcal{E}(\sigma))$ $=\frac{1}{d^{2}}\sum_{a,b,c}|\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}D_{a}\widetilde{P}^{\prime}_{b}E_{c})|^{2}$
$=\frac{1}{d^{2}}\sum_{a,b,c}|\operatorname{\operatorname{tr}}(D_{a}\widetilde{P}^{\prime}_{b}\widetilde{P}^{\prime}_{b}E_{c}\widetilde{P}_{\infty})|^{2}$
$\leq\frac{1}{d^{2}}\sum_{a,b,c}\operatorname{\operatorname{tr}}(D_{a}\widetilde{P}^{\prime}_{b}D_{a}^{*})\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}E_{c}^{*}\widetilde{P}^{\prime}_{b}E_{c})$
$\overset{(a)}{=}\frac{1}{d^{2}}\sum_{b}\Bigl(\operatorname{\operatorname{tr}}(\widetilde{P}^{\prime}_{b})\sum_{c}\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}E_{c}^{*}\widetilde{P}^{\prime}_{b}E_{c})\Bigr)$
$\leq\frac{\max_{b}\operatorname{\operatorname{tr}}(\widetilde{P}^{\prime}_{b})}{d^{2}}\sum_{b,c}\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}E_{c}^{*}\widetilde{P}^{\prime}_{b}E_{c})$
$\overset{(b)}{=}\frac{\max_{b}\operatorname{\operatorname{tr}}(\widetilde{P}^{\prime}_{b})}{d^{2}}\sum_{c}\operatorname{\operatorname{tr}}(\widetilde{P}_{\infty}E_{c}^{*}E_{c})$
[…]
$=\frac{\|\lambda(\mathcal{G})\|_{\infty}}{\|\lambda(\mathcal{F})\|_{\infty}}.$
For $(a)$, we used the fact that the operators $\{D_{a}\}_{a}$ are the Kraus operators of $\mathcal{F}_{1}\mathcal{D}\mathcal{G}_{2}$, which is a channel and hence trace preserving. For $(b)$, we used that $\sum_{b}\widetilde{P}^{\prime}_{b}=\mathds{1}_{\mathcal{G}_{1}}$. Finally, for $(c)$, we used the Kraus relation $\sum_{c}E^{*}_{c}E_{c}=\mathds{1}_{\mathcal{F}}$ for the channel $\mathcal{G}_{1}\circ\mathcal{E}$. This ends the proof of eq. (38). ∎
We can now present the proof of Theorem 2.
Proof of Theorem 2.
Let $\mathcal{F}$, $\mathcal{G}$ be two idempotent channels, $(k_{\nu})_{\nu\in\mathbb{N}}$, $(n_{\nu})_{\nu\in\mathbb{N}}$ be two integer sequences and $\operatorname{\varepsilon}>0$ satisfying the assumptions of Theorem 2. Let $p_{\min}\in\{1,+\infty\}$ such that
$\min_{p\in\{1,+\infty\}}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}=\frac{\log(\|\lambda(\mathcal{G})\|_{p_{\min}})}{\log(\|\lambda(\mathcal{F})\|_{p_{\min}})}.$
Then, we have
$\lim_{\nu\to\infty}\frac{k_{\nu}}{n_{\nu}}\geq\frac{\log(\|\lambda(\mathcal{G})\|_{p_{\min}})}{\log(\|\lambda(\mathcal{F})\|_{p_{\min}})}+\operatorname{\varepsilon}.$
Then, for all $\nu$ big enough,
$\frac{k_{\nu}}{n_{\nu}}\geq\frac{\log(\|\lambda(\mathcal{G})\|_{p_{\min}})}{\log(\|\lambda(\mathcal{F})\|_{p_{\min}})}+\frac{\operatorname{\varepsilon}}{2},$
so that
$\frac{1}{\|\lambda(\mathcal{F})\|^{\frac{\operatorname{\varepsilon}}{2}n_{\nu}}_{p_{\min}}}\geq\frac{\|\lambda(\mathcal{G})\|^{n_{\nu}}_{p_{\min}}}{\|\lambda(\mathcal{F})\|^{k_{\nu}}_{p_{\min}}}.$
As, by assumption, $n_{\nu}\underset{\nu\to\infty}{\to}\infty$ and $\|\lambda(\mathcal{F})\|_{p_{\min}}>1$, we get that
$\lim_{\nu\to\infty}\frac{\|\lambda(\mathcal{G})\|^{n_{\nu}}_{p_{\min}}}{\|\lambda(\mathcal{F})\|^{k_{\nu}}_{p_{\min}}}=0.$ (46)
However, by Theorem 8, for all $\nu\in\mathbb{N}$, all encoding and decoding channels $\mathcal{E}_{\nu}$, $\mathcal{D}_{\nu}$,
$\frac{1}{2}\|\mathcal{F}^{\otimes k_{\nu}}-\mathcal{D}_{\nu}\mathcal{G}^{\otimes n_{\nu}}\mathcal{E}_{\nu}\|_{\diamond}\geq 1-\frac{\|\lambda(\mathcal{G}^{\otimes n_{\nu}})\|_{p_{\min}}}{\|\lambda(\mathcal{F}^{\otimes k_{\nu}})\|_{p_{\min}}}=1-\frac{\|\lambda(\mathcal{G})\|^{n_{\nu}}_{p_{\min}}}{\|\lambda(\mathcal{F})\|^{k_{\nu}}_{p_{\min}}},$
where we used for the right-hand side equality the fact that both the shape vector and the $\ell_{p}$-norms are compatible with the tensor product, i.e. Proposition 4. Then, Theorem 2 follows by injecting eq. (46) in this last inequality. ∎
Acknowledgements
We thank Robert Salzmann for helpful discussions about [9] and [10], as well as about sets of approximate fixed point of channels. ID and OF acknowledge funding by the European Research Council (ERC Grant AlgoQIP, Agreement No. 851716). LG acknowledges funding by National Natural Science Foundation of China (NNFSC) via grant No.12401163. This work started when MR was a Marie Skłodowska-Curie Fellow at ENS de Lyon, and he acknowledges funding from the European Union’s Horizon Research and Innovation programme, grant Agreement No. HORIZON-MSCA-2022-PF-01 (Project number: 101108117).
Appendix A An application of approximate $C^{*}$-algebras
In a recent article, Kitaev introduced and studied approximate $C^{*}$-algebras and almost idempotent quantum channels, in order to find an algebraic framework which allows to naturally express the effects of the noise on quantum information [9]. As most of our proofs in Section 4 exploit the decomposition properties of finite-dimensional *-algebras, we use approximate $C^{*}$-algebras to strengthen the converse of Theorem 1 and show that it holds even if a small error is allowed.
Theorem 10.
[…]
$\|\mathcal{F}^{\otimes k}-\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}\|_{\diamond}\leq\frac{\delta_{\max}\lambda_{\min,n}}{6d^{k}},$
then
$\frac{k}{n}\leq\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}.$
Note that in Theorem 10, the error allowed depends on the dimension of the input space of $\mathcal{F}$, on the number of copies of $\mathcal{F}$ obtained and the number of copies of $\mathcal{G}$ used as well as on the encoder used for the emulation. This theorem is therefore less operationally motivated than Theorems 1 and 2, but it still strengthens the converse of Theorem 1 and its proof leads to nice mathematical developments, especially on multiplicative domains (see Theorem 12).
Note that an interesting direction would be to extend our Theorem 1 to the problem of emulating almost idempotent channels one with another, but we leave this for future work.
The Appendix is divided as follows. In Subsection A.1, we recall, for completeness, the Error Reduction Theorem for approximate $C^{*}$-algebras. We then move on to proving Theorem 12 on approximate multiplicative domain of maps satisfying the Kadison-Schwarz inequality in Subsection A.2, which will be used in the proof of Theorem 10 in Subsection A.3.
Note that this Appendix assumes familiarity with basic concepts from operator algebras.
A.1 The Error Reduction Theorem
We begin by recalling the definitions of $\operatorname{\varepsilon}$-$C^{*}$-algebras and $\delta$-homomorphisms.
Definition 1 ($\operatorname{\varepsilon}$-$C^{*}$-algebra, Definition 2.1 in [9]).
An $\operatorname{\varepsilon}$-Banach algebra is a Banach space $\mathcal{A}$, endowed with a bilinear multiplication map $\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A}$ such that
$\forall x,y\in\mathcal{A},~~\|xy\|\leq(1+\operatorname{\varepsilon})\|x\|\|y\|,$
$\forall x,y,z\in\mathcal{A},~~\|(xy)z-x(yz)\|\leq\operatorname{\varepsilon}\|x\|\|y\|\|z\|.$
A $\operatorname{\varepsilon}-$Banach algebra is a complex $\operatorname{\varepsilon}$-Banach algebra with a conjugate linear involution $x\mapsto x^{*}$ satisfying the equations:
$\forall x,y\in\mathcal{A},~\|x^{*}\|=\|x\|,~~(xy)^{*}=y^{*}x^{*}.$ (47)
An $\operatorname{\varepsilon}$-$C^{*}$-algebra is one satisfying the following property:
$\forall x\in\mathcal{A},~\|x^{*}x\|\geq(1-\operatorname{\varepsilon})\|x\|^{2}.$ (48)
The unit element $\mathds{1}\in\mathcal{A}$ should satisfy the following conditions:
$\|x\mathds{1}-x\|\leq\operatorname{\varepsilon}\|x\|,~\|\mathds{1}x-x\|\leq\operatorname{\varepsilon}\|x\|,~|\|\mathds{1}\|-1|\leq\operatorname{\varepsilon}.$ (49)
If $\mathcal{A}$ is involutive, we also have
$\mathds{1}^{*}=\mathds{1}.$ (50)
A $\delta$-homomorphism is then naturally defined in the following way.
Definition 2 ($\delta$-homomorphism, Definition 2.2 in [9]).
A $\delta$-homomorphism from an $\operatorname{\varepsilon}^{\prime}$-Banach algebra $\mathcal{A}^{\prime}$ to an $\operatorname{\varepsilon}^{\prime\prime}$-Banach algebra $\mathcal{A}^{\prime\prime}$ is a bounded linear map $v:\mathcal{A}^{\prime}\rightarrow\mathcal{A}^{\prime\prime}$ that almost preserves the unit and the multiplication:
$\forall x,y\in\mathcal{A}^{\prime},~\|v(\mathds{1})-\mathds{1}\|\leq\delta,$ (51)
$\|v(xy)-v(x)v(y)\|\leq\delta\|x\|\|y\|.$
A non-unital $\delta$-homomorphism is defined by imposing only the second condition. In the *-algebra setting, it is also required that $v(x^{*})=v(x)^{*}$.
A $\delta$-inclusion is a $\delta$-homomorphism such that
$\forall x\in\mathcal{A}^{\prime},~(1-\delta)\|x\|\leq\|v(x)\|\leq(1+\delta)\|x\|.$
The Error Reduction Theorem then states that if there is a $\delta$-inclusion of an exact $C^{*}$-algebra $\mathcal{A}$ into an $\operatorname{\varepsilon}$-$C^{*}$-algebra $\mathcal{B}$ for $\delta\leq\delta_{\max}$ and $\operatorname{\varepsilon}\leq\operatorname{\varepsilon}_{\max}$, which are constants that do not depend on the dimensions of $\mathcal{A}$ and $\mathcal{B}$, one can ‘correct’ this $\delta$-inclusion into a less noisy approximate inclusion.
Theorem 11 (Error Reduction Theorem, Corollary 8.3 in [9]).
There exists some positive constants $\operatorname{\varepsilon}_{max}$, $\delta_{max}$ and $c_{0}$ such that for all $\operatorname{\varepsilon}<\operatorname{\varepsilon}_{max}$, if a finite-dimensional $C^{*}$-algebra $\mathcal{A}$ is $\delta_{max}$-included into an $\operatorname{\varepsilon}$-$C^{*}$-algebra $\mathcal{B}$, then there is also a $(c_{0}\operatorname{\varepsilon})$-inclusion. If the original inclusion is bijective, then so is the new inclusion.
Note that as every $C^{*}$-algebra is a 0-$C^{*}$-algebra, Theorem 11 states that when we have a $\delta_{\max}$-inclusion between two $C^{*}$-algebras, it is possible to lift it to an exact inclusion. This can be interpreted as a case of complete correction in the sense that the reduction of the error parameter $\delta$ is total.
A.2 Approximate Multiplicative Domains
[…]
$\{x\in\mathcal{L}(\mathcal{H}^{\prime})~|~\mathcal{F}^{*}(x^{*}x)=\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(x)\text{ and }\mathcal{F}^{*}(xx^{*})=\mathcal{F}^{*}(x)\mathcal{F}^{*}(x^{*})\}.$
[…]
$\mathcal{C}_{\delta}^{R}$ $=\{x\in\mathcal{L}(\mathcal{H}^{\prime})|\|\mathcal{F}^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(x)\|\leq\delta\|x\|^{2}\},$
$\mathcal{C}_{\delta}^{L}$ $=\{x\in\mathcal{L}(\mathcal{H}^{\prime})|\|\mathcal{F}^{*}(xx^{*})-\mathcal{F}^{*}(x)\mathcal{F}^{*}(x^{*})\|\leq\delta\|x\|^{2}\},$
$\mathcal{C}_{\delta}$ $=\mathcal{C}_{\delta}^{R}\cap\mathcal{C}_{\delta}^{L}.$
We can show the following theorem.
Theorem 12.
[…]
$\forall x,y\in\mathcal{C}_{\delta},\|\mathcal{F}^{*}(xy)-\mathcal{F}^{*}(x)\mathcal{F}^{*}(y)\|\leq 2\delta\|x\|\|y\|.$ (52)
More precisely, we have that
$\forall x\in\mathcal{C}^{L}_{\delta},\forall y\in\mathcal{C}^{R}_{\delta},\|\mathcal{F}^{*}(xy)-\mathcal{F}^{*}(x)\mathcal{F}^{*}(y)\|\leq 2\delta\|x\|\|y\|.$ (53)
The proof of this theorem uses the following lemma.
Lemma 13.
Let $\{x_{k}\in\mathcal{L}(\mathcal{H})~|~k\in[K]\}$ and $\{\widetilde{x}_{k}\in\mathcal{L}(\mathcal{H})~|~k\in[K]\}$ be two sets of operators such that for all $k$ even,
$x_{k}\leq\widetilde{x}_{k},$
and for $k$ odd,
$x_{k}=\widetilde{x}_{k}.$
Then, for all $t\in\mathbb{R}$,
$\sum_{k=1}^{K}t^{k}x_{k}\leq\sum_{k=1}^{K}t^{k}\widetilde{x}_{k}.$ (54)
Proof of Lemma 13.
Let $t\in\mathbb{R}$:
$\sum_{k=1}^{K}t^{k}\widetilde{x}_{k}-\sum_{k=1}^{K}t^{k}x_{k}=\sum_{k\in[K],~k\text{ even}}t^{k}(\widetilde{x}_{k}-x_{k})\geq 0,$
where the right-hand side inequality follows from the assumption that $\widetilde{x}_{k}\geq x_{k}$ for all $k$ even and also that, for all $t\in{\mathbb{R}}$, $t^{k}\geq 0$ for $k$ even. ∎
Proof of Theorem 12.
This proof generalises the ones of Theorems 5.4 and 5.7 in [18]. In this proof, we let $\mathds{1}$ denote the identity operator of $\mathcal{L}(\mathcal{H}^{\prime})$.
We have that $x\in\mathcal{C}^{L}_{\delta}$ if and only if $x^{*}\in\mathcal{C}^{R}_{\delta}$ and vice versa. Therefore, it suffices to show that for all $x,y\in\mathcal{C}^{R}_{\delta}$,
$\|\mathcal{F}^{*}(x^{*}y)-\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(y)\|\leq 2\delta\|x\|\|y\|.$ (55)
Let $x,y\in\mathcal{C}^{R}_{\delta}$, $t\in{\mathbb{R}}$ and $z=tx+y$. As $\mathcal{F}^{*}$ satisfies the Kadison-Schwarz inequality,
$0\leq\mathcal{F}^{*}(z^{*}z)-\mathcal{F}^{*}(z^{*})\mathcal{F}^{*}(z).$
If we develop the right-hand side, we get:
$0\leq t^{2}Q+tV+W,$
with
$Q$ $=\mathcal{F}^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(x),$
$W$ $=\mathcal{F}^{*}(y^{*}y)-\mathcal{F}^{*}(y^{*})\mathcal{F}^{*}(y),$
$v$ $=\mathcal{F}^{*}(x^{*}y)-\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(y),$
$V$ $=v+v^{*}.$
Then, as $\mathcal{F}^{*}$ satisfies the Kadison-Schwarz inequality, $0\leq Q$ and $0\leq W$. Furthermore, $Q\leq\|Q\|\mathds{1}\leq\delta\|x\|^{2}\mathds{1}$ and $W\leq\|W\|\mathds{1}\leq\delta\|y\|^{2}\mathds{1}$ by assumption. Therefore, by Lemma 13, for all $t\in{\mathbb{R}}$:
$0\leq t^{2}Q+tV+W\leq t^{2}\delta\|x\|^{2}\mathds{1}+tV+\delta\|y\|^{2}\mathds{1}.$
As $V=v+v^{*}$ is Hermitian, we have $V=UDU^{*}$ with $U$ a unitary matrix and $D$ a diagonal matrix with real coefficients. Conjugating the last inequality by $U$, we get:
$0\leq t^{2}\delta\|x\|^{2}\mathds{1}+tD+\delta\|y\|^{2}\mathds{1}.$
The right-hand side of this equation is simply a diagonal operator, whose diagonal coefficients are quadratic polynomials in $t$. As this equation gives us that this operator is positive semidefinite for all $t\in{\mathbb{R}}$, each of these quadratic polynomials is always non-negative, and can thus have at most a single root. Therefore, if we write $\{D_{i}\}_{i}$ the diagonal entries of $D$, the discriminant of the polynomial
$\delta\|x\|^{2}t^{2}+D_{i}t+\delta\|y\|^{2},$
has to be non-positive, that is
$D_{i}^{2}\leq 4\delta^{2}\|x\|^{2}\|y\|^{2}.$
So that:
$|D_{i}|\leq 2\delta\|x\|\|y\|.$
As the Schatten $\infty$-norm of $V=UDU^{*}$ is $\|V\|=\max_{i}|D_{i}|$, we have:
$\|v+v^{*}\|=\|V\|\leq 2\delta\|x\|\|y\|.$
We can apply the same line of argument replacing $y$ with $iy$ to find
$\|iv-iv^{*}\|\leq 2\delta\|x\|\|y\|.$ (56)
We can decompose $v=h_{1}+ih_{2}$ with $h_{1}$ and $h_{2}$ Hermitian operators [^6] Taking $h_{1}=\frac{1}{2}(v+v^{*})$, $h_{2}=\frac{1}{2i}(v-v^{*})$., to find:
$\|v+v^{*}\|=2\|h_{1}\|\leq 2\delta\|x\|\|y\|,$
$\|iv-iv^{*}\|=2\|h_{2}\|\leq 2\delta\|x\|\|y\|.$
Therefore,
$\|v\|\leq\|h_{1}\|+\|h_{2}\|\leq 2\delta\|x\|\|y\|,$ (57)
which is what we claimed. ∎
A.3 Proof of Theorem 10
Finally, we prove in this subsection Theorem 10 in the same way as we proved the converse of Theorem 1, that is by first proving the one-shot case and then using the compatibility of the shape vector and the $\ell_{p}$-norms with the tensor product to conclude.
Proposition 7.
[…]
$\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\leq\frac{\delta_{\max}\lambda_{\min,1}}{6d},$ (58)
with $d=\dim(\mathcal{H}_{\mathcal{F}})$, $\delta_{\max}$ is the constant of Theorem 11 and $\lambda_{\min,1}$ is the minimal non-zero eigenvalue of $\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})$. We write $e_{\mathcal{G}\mathcal{E}}=\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})^{0}$ the orthogonal projector on the support $\operatorname{\operatorname{supp}}(\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}}))$ and we define
$\widetilde{\mathcal{G}}^{*}:x\mapsto e_{\mathcal{G}\mathcal{E}}\mathcal{G}^{*}(x)e_{\mathcal{G}\mathcal{E}}.$
Then, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\delta_{\max}$-inclusion from the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$ into the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$. By the Error Reduction Theorem 11, there is therefore an injective unital *-homomorphism embedding $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$, thus, for all $p\in[1,+\infty]$,
$\|\lambda(\mathcal{F})\|_{p}\leq\|\lambda(\mathcal{G})\|_{p}.$
We divide the proof that the map $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to $\operatorname{Rg}(\mathcal{F}^{*})$ satisfies all the properties of a $\delta_{\max}$-inclusion into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ in multiple lemmas, each of which being devoted to prove one of these properties.
[…]
$\|\mathcal{F}\|_{\text{cb}}=\sup_{n\in\mathbb{N}^{*}}\|\mathcal{F}\otimes\mathrm{Id}_{n}\|_{\infty\rightarrow\infty},$
with $\|\cdot\|_{\infty\rightarrow\infty}$ the ‘infinity-to-infinity’ norm, i.e. the map norm of linear maps from $\mathcal{L}(\mathcal{H})$ to $\mathcal{L}(\mathcal{H}^{\prime})$ when the input and output spaces are endowed with the Schatten $\infty$-norms. That is:
$\|\mathcal{F}\|_{\infty\to\infty}=\sup_{\|x\|\leq 1}\|\mathcal{F}(x)\|.$
[…]
$\|\mathcal{F}\|_{\diamond}=\|\mathcal{F}^{*}\|_{\text{cb}}.$
Lemma 14 (Almost norm preservation).
Let $\mathcal{F}$, $\mathcal{G}$, $\widetilde{\mathcal{G}}$, $\mathcal{E}$ and $\mathcal{D}$ be defined as in the statement of Proposition 7 and let
$\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\text{cb}}.$
Then, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
$(1-\delta)\|x\|\leq\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|\leq\|x\|.$
Note that this lemma implies that as long as $\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\leq\delta<1$, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is injective as every element of its kernel $x\in\ker(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*})$ satisfies $(1-\delta)\|x\|\leq\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|=0$, hence $\|x\|=0$ so that $x=0$.
Proof.
The right-hand side inequality follows from the fact that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}:\mathcal{L}(\mathcal{H}_{\mathcal{F}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ is a subunital completely positive map (when the identity considered in the output space of $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$). For the left-hand side inequality, recall that, by Lemma 7, for all $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}})$,
$\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x).$
Then, we have for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$:
$\|x\|-\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|$ $\overset{(a)}{\leq}\|x\|-\|\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|$
$\overset{(b)}{\leq}\|x-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}(x)\|$
$\overset{(c)}{=}\|\mathcal{F}^{*}(x)-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}(x)\|$
$\leq\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{cb}\|x\|$
$\leq\delta\|x\|,$
where $(a)$ follows from the fact that $\|\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|\leq\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|$, since $\mathcal{E}^{*}\mathcal{G}^{*}$ is a unital completely positive map. Then, $(b)$ follows from the triangular inequality. Finally, $(c)$ follows from the fact that $\mathcal{F}^{*}(x)=x$ for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$. The last inequality can be rewritten as
$(1-\delta)\|x\|\leq\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|.$
We then move on to the proof that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}:\operatorname{Rg}(\mathcal{F}^{*})\rightarrow e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is unital, where we recall that the unit of the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is $e_{\mathcal{G}\mathcal{E}}$, as, for every element $x=e_{\mathcal{G}\mathcal{E}}x^{\prime}e_{\mathcal{G}\mathcal{E}}\in e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$, $e_{\mathcal{G}\mathcal{E}}x=xe_{\mathcal{G}\mathcal{E}}=x$. Also note that for a general $\delta$-inclusion, one only requires almost unitality.
Lemma 15 (Exact unitality).
Let $\mathcal{F}$, $\mathcal{G}$, $\widetilde{\mathcal{G}}$, $\mathcal{E}$ and $\mathcal{D}$ be defined as in Proposition 7. Then:
$\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=e_{\mathcal{G}\mathcal{E}}.$
Proof.
As $\mathcal{G}$ and $\mathcal{D}$ are assumed to be channels, their adjoints are unital maps on $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$, so that
$\mathcal{G}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=\mathds{1}_{\mathcal{G}}.$
Therefore
$\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=e_{\mathcal{G}\mathcal{E}}\mathcal{G}^{*}(\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}}))e_{\mathcal{G}\mathcal{E}}=e_{\mathcal{G}\mathcal{E}}\mathds{1}_{\mathcal{G}}e_{\mathcal{G}\mathcal{E}}=e_{\mathcal{G}\mathcal{E}}.$
Finally, to prove the almost preservation of the multiplication on $\operatorname{Rg}(\mathcal{F}^{*})$ by $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$, we will need the following additional lemma.
Lemma 16.
Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be a channel, write $d=\dim(\mathcal{H})$, $e=\Phi(\mathds{1})^{0}$ and $\lambda_{\min,1}$ the minimal non-zero eigenvalue of $\Phi(\mathds{1})$. Then, for all positive semidefinite $x\in\mathcal{L}(\mathcal{H}^{\prime})$,
$\frac{\lambda_{\min,1}}{d}\|exe\|\leq\|\Phi^{*}(x)\|.$ (59)
Proof.
Let $x\in\mathcal{L}(\mathcal{H}^{\prime})$ be positive semidefinite, then
$\|\Phi^{*}(x)\|$ $\geq\frac{1}{d}\operatorname{\operatorname{tr}}(\Phi^{*}(x))$
$=\frac{1}{d}\operatorname{\operatorname{tr}}(x\Phi(\mathds{1}))$
$\geq\frac{\lambda_{\min,1}}{d}\operatorname{\operatorname{tr}}(exe)$
$\geq\frac{\lambda_{\min,1}}{d}\|exe\|.$
Note that this lemma implies the second statement of Lemma 7. We can now move on to the proof of the almost preservation of the multiplication, which uses our Theorem 12 on the approximate multiplicative domain of maps satisfying the Kadison-Schwarz inequality.
Lemma 17 (Almost preservation of the multiplication).
Let $\mathcal{F}$, $\mathcal{G}$, $\widetilde{\mathcal{G}}$, $\mathcal{E}$ and $\mathcal{D}$ be defined as in the statement of Proposition 7, let
$\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond},$
let $\lambda_{\min,1}$ be the minimal non-zero eigenvalue of $\widetilde{\mathcal{G}}^{*}\mathcal{E}^{*}(\mathds{1}_{\mathcal{F}})$, $d=\dim(\mathcal{H}_{\mathcal{F}})$ and let $x$, $y\in\operatorname{Rg}(\mathcal{F}^{*})$. We then have:
$\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(xy)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(y)\|\leq\frac{6\delta d}{\lambda_{\min,1}}\|x\|\|y\|.$
Proof.
We prove that, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
$\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|$ $\leq\frac{3\delta d}{\lambda_{\min,1}}\|x\|^{2},$
$\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(xx^{*})-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\|$ $\leq\frac{3\delta d}{\lambda_{\min,1}}\|x\|^{2},$
and then use Theorem 12 to conclude. First, to simplify our notations, we let $\Phi^{*}=\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$. We have:
$\|\Phi^{*}(x^{*}x)-\Phi^{*}(x^{*})\Phi^{*}(x)\|$ $=\|\Phi^{*}(x^{*}x)-\Phi^{*}(x^{*})\Phi^{*}(x)+x^{*}x-x^{*}x\|$
$\leq\|\Phi^{*}(x^{*}x)-x^{*}x\|+\|x^{*}x-\Phi^{*}(x^{*})\Phi^{*}(x)+x^{*}\Phi^{*}(x)-x^{*}\Phi^{*}(x)\|$
$\overset{(a)}{\leq}\delta\|x\|^{2}+\|x\|\|x-\Phi^{*}(x)\|+\|x^{*}-\Phi^{*}(x^{*})\|\|\Phi^{*}(x)\|$
$\overset{(b)}{\leq}3\delta\|x\|^{2},$
where $(a)$ follows from the fact that as $x\in\operatorname{Rg}(\mathcal{F}^{*})$, we have
$\|\Phi(x^{*}x)-x^{*}x\|=\|\Phi^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*}x)\|\leq\|\Phi^{*}-\mathcal{F}^{*}\|_{\text{cb}}\|x^{*}x\|\leq\delta\|x\|^{2}.$
By the same argument, we also have that $\|x-\Phi^{*}(x)\|\leq\delta\|x\|$ and $\|x^{*}-\Phi^{*}(x^{*})\|\leq\delta\|x\|$, hence $(b)$ follows.
Then, applying the Kadison-Schwarz inequality, we get:
$\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)\geq\mathcal{E}^{*}\mathcal{G}^{*}(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x))\geq\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x).$
Therefore:
$\|\mathcal{E}^{*}\mathcal{G}^{*}(\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x))\|\leq\|\Phi^{*}(x^{*}x)-\Phi^{*}(x^{*})\Phi^{*}(x)\|\leq 3\delta\|x\|^{2}.$
Thus, by Lemma 16, we obtain
$\|\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*}x)-\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x^{*})\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)\|\leq\frac{3d\delta}{\lambda_{\min,1}}\|x\|^{2}.$
We can apply the exact same line of reasoning to prove the same inequality but with the roles of $x$ and $x^{*}$ swapped. ∎
Gathering all these lemmas, the proof of Proposition 7 is then as follows.
Proof of Proposition 7.
Let $\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}.$ By Lemmas 1415 and 17, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\max\{\delta,\frac{6d\delta}{\lambda_{\min,1}}\}$-inclusion (we trivially have that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is Hermitian preserving). However, as
$d=\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{F}})=\operatorname{\operatorname{tr}}(\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}}))\geq\lambda_{\min,1},$
$\frac{d}{\lambda_{\min,1}}\geq 1$ so that $\delta\leq\frac{6d}{\lambda_{\min,1}}\delta$. Finally, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\frac{6d\delta}{\lambda_{\min,1}}$-inclusion of $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$.
Then, if we suppose that
$\frac{6d\delta}{\lambda_{\min,1}}\leq\delta_{\max},$
i.e.
$\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\leq\delta\leq\frac{\delta_{\max}\lambda_{\min,1}}{6d},$
by the Error Reduction Theorem (Theorem 11), there is an injective *-homomorphism from $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$, thus by Kuperberg’s theorem (Theorem 6) and the second part of Proposition 6, for all $p\in[1,+\infty]$,
$\|\lambda(\mathcal{F})\|_{p}\leq\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\mathcal{G})\|_{p}.$
We can finally show Theorem 10.
Proof of Theorem 10.
Basically, we simply apply Proposition 7 to $\mathcal{F}^{\otimes k}:\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})$} and $\mathcal{G}^{\otimes n}:\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$. Just note that the dimension of $\mathcal{H}_{\mathcal{F}}^{\otimes k}$ is $d^{k}$, with $d=\dim(\mathcal{H}_{\mathcal{F}})$. Thus, by Proposition 7, if the hypotheses of Theorem 10 are met, for all $p\in[1,+\infty]$, we have:
$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}\leq\|\lambda(\mathcal{G}^{\otimes n})\|_{p}.$
This imply, by the multiplicativity of the $\ell_{p}$-norms of the shape vector with the tensor product, that
$\frac{k}{n}\leq\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}=C(\mathcal{G}\mapsto\mathcal{F}),$
where the right-hand side equality is the first statement in Theorem 1.

### v2

@@
 We start by considering the zero-error regime.
-We say that a channel $\mathcal{G}$ can emulate another channel $\mathcal{F}$ with rate $r$, if there exist an encoding and a decoding channel $\mathcal{E}$ and $\mathcal{D}$ (respectively) as well as a pair of integers $(k,n)$ such that
+We say that a channel $\mathcal{G}$ can emulate another channel $\mathcal{F}$ with rate $r$, if there exists an encoding and a decoding channel $\mathcal{E}$ and $\mathcal{D}$ (respectively) as well as a pair of integers $(k,n)$ such that
 $\mathcal{F}^{\otimes k}=\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E},$
@@
 Theorem 1.
-Let $\mathcal{F}=\mathcal{F}\circ\mathcal{F}$ and $\mathcal{G}=\mathcal{G}\circ\mathcal{G}$ be two idempotent channels with $\lambda(\mathcal{F})\neq(1)$ [^1] If $\lambda(\mathcal{F})=(1)$, then $\mathcal{F}$ is a replacer channel (see Proposition 1) and it is simple to see that we can emulate it with any other idempotent channel $\mathcal{G}$ with an arbitrarily large rate., then
+Let $\mathcal{F}=\mathcal{F}\circ\mathcal{F}$ and $\mathcal{G}=\mathcal{G}\circ\mathcal{G}$ be two idempotent channels with $\lambda(\mathcal{F})\neq(1)$ [^1] If $\lambda(\mathcal{F})=(1)$, then $\mathcal{F}$ is basically a replacer channel (see Proposition 1) and it is simple to see that we can emulate it with any other idempotent channel $\mathcal{G}$ with an arbitrarily large rate., then
 $C(\mathcal{G}\mapsto\mathcal{F})=\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}.$ (1)
@@
 This follows from the fact that the infimum in the formulas for $C(\mathcal{G}\mapsto\mathcal{F})$ and $C(\mathcal{F}\mapsto\mathcal{G})$ can be reached for different values of $p$.
-Next, we consider the setting beyond the zero-error regime, where the emulation does not need to be exact: the rate $r$ is achievable with error $\delta\in[0,1]$ if there exists channel $\mathcal{E},\mathcal{D}$ and $k,n$ with $r\leq k/n$ such that
+Next, we consider the setting beyond the zero-error regime, where the emulation does not need to be exact: the rate $r$ is achievable with error $\delta\in[0,1]$ if there exists channels $\mathcal{E},\mathcal{D}$ and $k,n$ with $r\leq k/n$ such that
 $\frac{1}{2}\|\mathcal{F}^{\otimes k}-\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}\|_{\diamond}\leq\delta,$
@@
 Finally, we end this section by introducing the notions of finite-dimensional *-algebras and their morphisms.
-A *-algebra $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H})$ is a $\mathbb{C}$-linear subspace of operators which is stable under the multiplication and and the * operation of $\mathcal{L}(\mathcal{H})$, i.e.
+In this article, we actually consider only *-subalgebras of $\mathcal{L}(\mathcal{H})$.
+Such a *-subalgebra $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H})$ is a linear subspace of $\mathcal{L}(\mathcal{H})$ which is stable under the multiplication and the * operation of $\mathcal{L}(\mathcal{H})$, i.e.
 taking the transpose conjugate.
@@
 $\mathcal{D}(\mathcal{H})$ Set of density operators in $\mathcal{L}(\mathcal{H})$
-$\mathcal{A},\mathcal{B}$ *-algebras
+$\mathcal{A},\mathcal{B}$ *-subalgebras of $\mathcal{L}(\mathcal{H})$
 $\rho$, $\sigma$ Density operators
@@
 $x^{0}$ Orthogonal projector on the support of $x$
-$x^{*}$ Hermitian conjugate of $x$
-$\|\cdot\|_{p}$ $\ell_{p}$-norm Ref. eq. (5)
+$x^{*}$ Hermitian adjoint of $x$
+$\|\cdot\|_{p}$ $\ell_{p}$-norm Eq. (5)
 $\mathcal{F}^{*}$ Adjoint of $\mathcal{F}$ for the Hilbert-Schmidt product
-$\operatorname{Rg}(\mathcal{F})$ Range of $\mathcal{F}$ Ref. eq. (6)
-$\lambda(\mathcal{F})$ Shape vector of $\mathcal{F}$ Ref. Prop. 1
+$\widehat{\mathcal{F}}$ Reduced channel associated to $\mathcal{F}$ Prop. 1, eq. (9)
+$\operatorname{Rg}(\mathcal{F})$ Range of $\mathcal{F}$
+$\lambda(\mathcal{F})$ Shape vector of $\mathcal{F}$ Prop. 1
 $[n]$ The set of integers $\{1,\dots,n\}$
@@
 In this section, we discuss the decomposition properties of the range of idempotent channels, which are summarized in the following propositions.
-The range of a map $f:\Omega\rightarrow\Omega^{\prime}$, denoted $\operatorname{Rg}(f)$, is the set of all possible images over the domain $\Omega$:
-$\operatorname{Rg}(f)=\{f(x)~|~x\in\Omega\}.$ (6)
-Proposition 1 (Structure of the range of idempotent channels).
+The range of a map $f:\Omega\rightarrow\Omega^{\prime}$, denoted $\operatorname{Rg}(f)$, is the set of all possible images over its domain $\Omega$, i.e.
+$\operatorname{Rg}(f)=\{f(x)~|~x\in\Omega\}$.
+The range of an idempotent channel coincides with its set of fixed points and the structure of this space is well understood, see for example the lecture notes [19].
+The following proposition summarizes the properties that we use.
+Proposition 1 (Structure of the range of idempotent channels, see [19]).
 […]
-$\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$ (7)
-and a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$, such that for all $x\in\mathcal{L}(\mathcal{H})$:
-$\mathcal{F}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}},$ (8)
+$\mathcal{H}=\mathcal{H}_{0}\oplus\mathcal{H}_{1},$ (6)
+with $\mathcal{H}_{0}=\operatorname{\operatorname{supp}}(\mathcal{F}(\mathds{1}))$ and $\mathcal{H}_{1}=\ker(\mathcal{F}(\mathds{1}))$, and $\mathcal{H}_{0}$ can further be decomposed as the orthogonal direct sum
+$\mathcal{H}_{0}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$ (7)
+so that there is a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$, such that
+$\operatorname{Rg}(\mathcal{F})=\Bigl(\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}\Bigr)\oplus 0.$ (8)
+Let $e=\mathcal{F}(\mathds{1})^{0}\in\mathcal{L}(\mathcal{H})$ be the orthogonal projector on $\mathcal{H}_{0}$, and $V:\mathcal{H}_{0}\rightarrow\mathcal{H}$ be the isometry embedding $\mathcal{H}_{0}$ into $\mathcal{H}$ such that $VV^{*}=e$ and $V^{*}V=\mathds{1}_{0}$ the identity operator in $\mathcal{L}(\mathcal{H}_{0})$.
+The reduced quantum channel associated to $\mathcal{F}$, denoted $\widehat{\mathcal{F}}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{0})$ is then defined as
+$\widehat{\mathcal{F}}:x\mapsto V^{*}\mathcal{F}(VxV^{*})V,$ (9)
+which is idempotent and whose range can be written as
+$\operatorname{Rg}(\widehat{\mathcal{F}})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$ (10)
+Furthermore, we can express $\widehat{\mathcal{F}}$ and its adjoint $\widehat{\mathcal{F}}^{*}$ in the following way:
+$\widehat{\mathcal{F}}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k},$ (11)
 and
-$\mathcal{F}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k},$ (9)
-where, for all $k\in[K]$, $P_{k}$ is the orthogonal projector onto $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, $d_{k}=\dim(\mathcal{H}_{k,1})\geq 1$, $m_{k}=\dim(\mathcal{H}_{k,2})$ and $\operatorname{\operatorname{tr}}_{k,2}:\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})\rightarrow\mathcal{L}(\mathcal{H}_{k,1})$ is the partial trace over $\mathcal{H}_{k,2}$, where the argument of this partial trace is seen as an operator in $\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})$ by restriction of its input and output to $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, which is a subspace of $\mathcal{H}$ for every $k\in[K]$.
-Furthermore, the range of $\mathcal{F}^{*}$ is a *-algebra and the following decompositions hold:
-$\operatorname{Rg}(\mathcal{F}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}},$ (10)
-$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$ (11)
-Note that $\mathcal{F}^{*}$ may not be faithful, indeed the states $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$ are not assumed to be full-rank.
+$\widehat{\mathcal{F}}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}},$ (12)
+where, for all $k\in[K]$, $P_{k}$ is the orthogonal projector onto $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, $d_{k}=\dim(\mathcal{H}_{k,1})\geq 1$, $m_{k}=\dim(\mathcal{H}_{k,2})$ and $\operatorname{\operatorname{tr}}_{k,2}:\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})\rightarrow\mathcal{L}(\mathcal{H}_{k,1})$ is the partial trace over $\mathcal{H}_{k,2}$.
+Therefore, the range of $\widehat{\mathcal{F}}^{*}$ is a unital *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$ and the following decomposition holds:
+$\operatorname{Rg}(\widehat{\mathcal{F}}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}.$ (13)
 We call the shape of $\mathcal{F}$ the $K$-dimensional vector whose coordinates are the dimensions $(d_{k})_{k\in[K]}$ sorted in non-increasing order, which we write $\lambda(\mathcal{F})$.
+[…]
 As an example of a shape vector, consider the qubit channel $\mathcal{F}:\mathcal{L}({\mathbb{C}}^{2})\rightarrow\mathcal{L}({\mathbb{C}}^{2})$ defined as
 $\mathcal{F}:x\mapsto\operatorname{\operatorname{tr}}(x)|0\rangle\!\langle 0|=\begin{pmatrix}\operatorname{\operatorname{tr}}(x)&0\\ 0&0\end{pmatrix}$,
-where the matrix expression is written for the canonical base of ${\mathbb{C}}^{2}$.
-Then, in the notations of eq.
-(7), we have $\mathcal{H}={\mathbb{C}}^{2}$, $K=1$, $\mathcal{H}_{k,1}={\mathbb{C}}$ and $\mathcal{H}_{k,2}={\mathbb{C}}^{2}$ so that $\lambda(\mathcal{F})=(1)$.
-Note also that, in this case, $\mathcal{F}$ is not faithful as $|0\rangle\!\langle 0|$ is not full-rank in $\mathcal{L}({\mathbb{C}}^{2})$.
-We remark that we have $\lambda(\mathcal{F})=(1)$ if and only if $\mathcal{F}$ is a replacer channel, i.e.
-$\mathcal{F}:x\mapsto\operatorname{\operatorname{tr}}(x)\rho$ with $\rho$ a fixed state.
+[…]
+$\mathcal{F}(\mathds{1})=\begin{pmatrix}2&0\\ 0&0\end{pmatrix}$.
+[…]
+$VV^{*}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=\mathcal{F}(\mathds{1})^{0}.$
+[…]
+$\widehat{\mathcal{F}}(x)=V^{*}\mathcal{F}(VxV^{*})V=x,$
+so that $\widehat{\mathcal{F}}$ is the identity on $\mathcal{L}({\mathbb{C}})$.
+Furthermore, we remark that we have $\lambda(\mathcal{F})=(1)$ if and only if its reduced channel is a replacer channel, i.e.
+$\widehat{\mathcal{F}}:x\mapsto\operatorname{\operatorname{tr}}(x)\rho$ with $\rho$ a fixed state.
 This fact follows easily from eq.
 (11) taking $K=1$.
+Before we present the proof of Proposition 1, we need to state the two following lemmas which will be used in the proof.
+Lemma 3.
+Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H})$ be a positive linear map [^2] Here ’positive’ means that $x\geq 0$ implies $\Phi(x)\geq 0$., let $\mathcal{H}_{0}=\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))$ and let $V:\mathcal{H}_{0}\rightarrow\mathcal{H}$ be an isometry embedding $\mathcal{H}_{0}$ into $\mathcal{H}$ so that $VV^{*}\in\mathcal{L}(\mathcal{H})$ is the orthogonal projection on $\mathcal{H}_{0}$.
+Then for all $x\in\mathcal{L}(\mathcal{H})$,
+$VV^{*}\Phi(x)VV^{*}=\Phi(x).$
+Proof.
+Let $x\in\mathcal{L}(\mathcal{H})$ be positive semidefinite, then
+$0\leq x\leq\|x\|\mathds{1},$
+so that by positivity of $\Phi$, we have
+$0\leq\Phi(x)\leq\|x\|\Phi(\mathds{1}).$
+Hence, $\Phi(x)$ is supported on the support of $\Phi(\mathds{1})$ which is, by definition, $\mathcal{H}_{0}$.
+Furthermore, as $x$ and $\mathds{1}$ are positive semidefinite, $\Phi(x)$ and $\Phi(\mathds{1})$ are positive semidefinite as well so that their respective range (as linear maps on $\mathcal{H}$) coincides with their support.
+Hence, we have
+$VV^{*}\Phi(x)VV^{*}=\Phi(x).$
+Now, if $x\in\mathcal{L}(\mathcal{H})$ is arbitrary, we can use the polarisation identity (see eq. (1.45) in [19]) to write
+$x=\frac{1}{4}\sum_{k=0}^{3}i^{k}(x+i^{k}\mathds{1})^{*}(x+i^{k}\mathds{1}),$
+so that $x$ is a linear combination of four positive semidefinite elements of $\mathcal{L}(\mathcal{H})$. Now, by linearity, we get
+$VV^{*}\Phi(x)VV^{*}=\Phi(x).$
+Furthermore, we remark that the set of fixed points and the range of an idempotent function coincide.
+Lemma 4.
+Let $f:\Omega\rightarrow\Omega$ be an arbitrary idempotent function, with $\Omega$ an arbitrary set, then
+$\operatorname{Rg}(f)=\{x\in\Omega~|~f(x)=x\}.$ (14)
+Proof.
+If $x$ is a fixed point, $x=f(x)\in\operatorname{Rg}(f)$. For the other inclusion, if $y=f(x)$, we also have $f(y)=f(f(x))=f(x)=y$, so $y$ is a fixed point. ∎
+Proposition 1 then basically follows from statements in Section 6.4 [19] and we detail it here only for the convenience of the reader. Similar derivation can be found in [1] in the Heisenberg picture.
+Proof of Proposition 1.
+We will first prove that the reduced channel $\widehat{\mathcal{F}}$, defined in eq.
+(9), is actually a channel.
+It is clearly completely positive as it is a composition of completely positive maps and it is trace preserving because for all $x\in\mathcal{L}(\mathcal{H}_{0})$,
+$\operatorname{\operatorname{tr}}(\widehat{\mathcal{F}}(x))=\operatorname{\operatorname{tr}}(V^{*}\mathcal{F}(VxV^{*})V)=\operatorname{\operatorname{tr}}(VV^{*}\mathcal{F}(VxV^{*})VV^{*})=\operatorname{\operatorname{tr}}(\mathcal{F}(VxV^{*}))=\operatorname{\operatorname{tr}}(VxV^{*})=\operatorname{\operatorname{tr}}(x),$
+where the third equality follows from Lemma 3. Furthermore, $\widehat{\mathcal{F}}$ is idempotent as, for $x\in\mathcal{L}(\mathcal{H}_{0})$,
+$\widehat{\mathcal{F}}(\widehat{\mathcal{F}}(x))$ $=V^{*}\mathcal{F}(VV^{*}\mathcal{F}(VxV^{*})VV^{*})V$
+$=V^{*}\mathcal{F}(\mathcal{F}(VxV^{*}))V$
+$=V^{*}\mathcal{F}(VxV^{*})V$
+$=\widehat{\mathcal{F}}(x).$
+The second equation follows from Lemma 3, and the third from the idempotence of $\mathcal{F}$.
+Furthermore, $\widehat{\mathcal{F}}$ admits a full-rank fixed point.
+Indeed, $\widehat{\mathcal{F}}(V^{*}\mathcal{F}(\mathds{1})V)$ is a fixed point of $\widehat{\mathcal{F}}$, since $\widehat{\mathcal{F}}$ is idempotent, and we have
+$\widehat{\mathcal{F}}(V^{*}\mathcal{F}(\mathds{1})V)=V^{*}\mathcal{F}(VV^{*}\mathcal{F}(\mathds{1})VV^{*})V=V^{*}\mathcal{F}\mathcal{F}(\mathds{1})V=V^{*}\mathcal{F}(\mathds{1})V,$
+which is full-rank in $\mathcal{H}_{0}=\operatorname{\operatorname{supp}}(\mathcal{F}(\mathds{1}))$, again using Lemma 3 for the second equality.
+Therefore, we can apply Theorem 6.12 in [19] to the adjoint $\widehat{\mathcal{F}}^{*}$ of $\widehat{\mathcal{F}}$.
+Hence, the set of fixed points of $\widehat{\mathcal{F}}^{*}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$.
+By Lemma 4, since $\widehat{\mathcal{F}}^{*}$ is idempotent, its set of fixed points coincides with its range, so that its range is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$.
+It is well known (see e.g.
+[17]), that if $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{0})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$, we can find a basis of $\mathcal{H}_{0}$ in which it can be expressed as an orthogonal direct sum as in eq.
+(13), which proves this equation.
+Furthermore, we get by the Proposition 1.5 of [19] that $\widehat{\mathcal{F}}^{*}$ takes the form of eq.
+(12).
+Then, taking the adjoint of this equation, we obtain that $\widehat{\mathcal{F}}$$\widehat{\mathcal{F}}^{**}$ can be written as in eq.
+(11).
+Then, for $x\in\mathcal{L}(\mathcal{H}_{0})$, we have that
+$\widehat{\mathcal{F}}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$
+Conversely, if
+$x=\sum_{k=1}^{K}x_{k}\otimes\rho_{k}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
+[…]
+Finally, to prove eq. (8), let $x\in\operatorname{Rg}(\mathcal{F})$. We then have
+$x=\mathcal{F}(x)=VV^{*}\mathcal{F}(x)VV^{*}=VV^{*}xVV^{*},$ (15)
+so that $x\in\mathcal{L}(\mathcal{H}_{0})\oplus 0$ where the 0 operator acts on $\mathcal{H}_{1}$. Then, $V^{*}xV\in\mathcal{L}(\mathcal{H}_{0})$ and
+$\widehat{\mathcal{F}}(V^{*}xV)=V^{*}\mathcal{F}(VV^{*}xVV^{*})V=V^{*}\mathcal{F}(x)V=V^{*}xV,$
+by Lemma 3 and by eq. (15). Hence,
+$V^{*}xV\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
+by eq. (10) and
+$x=VV^{*}xVV^{*}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}\oplus 0.$
+Conversely, if
+$x=\begin{pmatrix}x_{0}&0\\ 0&0\end{pmatrix}$,
+with
+$x_{0}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
+then
+$x=Vx_{0}V^{*}=V\widehat{\mathcal{F}}(x_{0})V^{*}=VV^{*}\mathcal{F}(Vx_{0}V^{*})VV^{*}=\mathcal{F}(Vx_{0}V^{*})=\mathcal{F}(x),$
+so that $x\in\operatorname{Rg}(\mathcal{F})$. ∎
+We will now show that if $\mathcal{F}$ and $\mathcal{G}$ are two idempotent channels, it is equivalent to study the emulation of $\mathcal{F}$ by $\mathcal{G}$ and the emulation of their reduced version, i.e.
+$\widehat{\mathcal{F}}$ by $\widehat{\mathcal{G}}$.
+First, we can explicitly find an encoder and a decoder channel to convert any idempotent channel $\mathcal{F}$ into its reduced version $\widehat{\mathcal{F}}$ and vice versa as shown in the following proposition.
+Proposition 2.
+[…]
+[…]
+Proof.
+Let $V:\mathcal{H}_{0}\rightarrow\mathcal{H}$ be the isometry embedding $\mathcal{H}_{0}$ into $\mathcal{H}$ which defines $\widehat{\mathcal{F}}$ as in Proposition 1.
+For the first equality, let $\widehat{\mathcal{E}}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}_{0})$ be defined by
+$\widehat{\mathcal{E}}:x\mapsto V^{*}\mathcal{F}(x)V,$
+and $\widehat{\mathcal{D}}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H})$ be defined by
+$\widehat{\mathcal{D}}:x\mapsto VxV^{*}.$
+Both $\widehat{\mathcal{E}}$ and $\widehat{\mathcal{D}}$ are completely positive trace preserving map.
+The complete positivity is easy.
+$\widehat{\mathcal{D}}$ is trace preserving because $\operatorname{\operatorname{tr}}(VxV^{*})=\operatorname{\operatorname{tr}}(V^{*}Vx)=\operatorname{\operatorname{tr}}(x)$ and $\widehat{\mathcal{E}}$ is trace preserving as well because for all $x\in\mathcal{L}(\mathcal{H})$,
+$\operatorname{\operatorname{tr}}(V^{*}\mathcal{F}(x)V)=\operatorname{\operatorname{tr}}(VV^{*}\mathcal{F}(x)VV^{*})=\operatorname{\operatorname{tr}}(\mathcal{F}(x))=\operatorname{\operatorname{tr}}(x),$
+where the second equality follows from Lemma 3. Then we have the factorisation
+$\mathcal{F}=\widehat{\mathcal{D}}\widehat{\mathcal{F}}\widehat{\mathcal{E}}$.
+Indeed, let $x\in\mathcal{L}(\mathcal{H})$, we have
+$\widehat{\mathcal{D}}\widehat{\mathcal{F}}\widehat{\mathcal{E}}(x)$ $=VV^{*}\mathcal{F}(VV^{*}\mathcal{F}(x)VV^{*})VV^{*}$
+$=\mathcal{F}(\mathcal{F}(x))$
+$=\mathcal{F}(x),$
+where the second equality follows from Lemma 3.
+Now, for the factorisation of $\widehat{\mathcal{F}}$, let $\mathcal{E}=\widehat{\mathcal{D}}$ and $\mathcal{D}=\widehat{\mathcal{E}}$, for $x\in\mathcal{L}(\mathcal{H}_{0})$, we have:
+$\mathcal{D}\mathcal{F}\mathcal{E}(x)$ $=\widehat{\mathcal{E}}\mathcal{F}\widehat{\mathcal{D}}(x)$
+$=V^{*}\mathcal{F}(\mathcal{F}(VxV^{*}))V$
+$=V^{*}\mathcal{F}(VxV^{*})V$
+$=\widehat{\mathcal{F}}(x).$
+As a consequence of this proposition, we get that studying the interconversion of two idempotent channels $\mathcal{F}$ and $\mathcal{G}$, is equivalent to study the interconversion of their reduced versions $\widehat{\mathcal{F}}$ and $\widehat{\mathcal{G}}$.
+Proposition 3.
+[…]
+$\inf_{\mathcal{D}\!,\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\inf_{\mathcal{D}\!,\mathcal{E}}\|\widehat{\mathcal{F}}-\mathcal{D}\widehat{\mathcal{G}}\mathcal{E}\|_{\diamond}$, (17)
+where the infimum are taken over encoders and decoders channels $\mathcal{E}$ and $\mathcal{D}$ between appropriate spaces.
+Proof.
+Let $\mathcal{E}:\mathcal{L}(\mathcal{H}_{\mathcal{F}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ and $\mathcal{D}:\mathcal{L}(\mathcal{H}_{\mathcal{G}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ be a pair of channels.
+Let $\mathcal{E}_{1}$, $\mathcal{D}_{1}$, $\mathcal{E}_{2}$, $\mathcal{D}_{2}$ be channels such that
+[…]
+which exist by Proposition 2. Then,
+$\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}$ $=\|\mathcal{F}-\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\|_{\diamond}$
+$=\|\mathcal{D}_{1}\|_{\diamond}\|\mathcal{F}-\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\|_{\diamond}\|\mathcal{E}_{1}\|_{\diamond}$
+$\geq\|\mathcal{D}_{1}\mathcal{F}\mathcal{E}_{1}-\mathcal{D}_{1}\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\mathcal{E}_{1}\|_{\diamond}$
+$=\|\widehat{\mathcal{F}}-\mathcal{D}_{1}\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\mathcal{E}_{1}\|_{\diamond},$
+where we used in the second equality that as $\mathcal{E}_{1}$, $\mathcal{D}_{1}$ are channels, $\|\mathcal{E}_{1}\|_{\diamond}=\|\mathcal{D}_{1}\|_{\diamond}=1$.
+Therefore,
+$\inf_{\mathcal{D}\!,\mathcal{E}}\|\widehat{\mathcal{F}}-\mathcal{D}\widehat{\mathcal{G}}\mathcal{E}\|_{\diamond}\leq\inf_{\mathcal{D}\!,\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}.$
+We can apply the same line of argument but converting $\widehat{\mathcal{F}}$ to $\mathcal{F}$ and $\mathcal{G}$ to $\widehat{\mathcal{G}}$ to prove the converse inequality. ∎
 […]
@@
 […]
-[…]
+We now show that the range and the shape vector of idempotent channels preserve the tensor product, a property that will be used in the proofs of our Theorems 1 and 2.
+Indeed, it will allow us to reduce the study of the problem of emulating $k$ copies of $\mathcal{F}$ using $n$ copies of $\mathcal{G}$ for general pairs of integers $(k,n)$ to the case where $k=n$.
 Proposition 5.
-[…]
-$\mathcal{F}=\mathcal{D}\circ\widetilde{\mathcal{F}}\circ\mathcal{E}$. (12)
+Let $\mathcal{F}_{1}:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{1})$ and $\mathcal{F}_{2}:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{L}(\mathcal{H}_{2})$ be two idempotent channels, we have
+$\operatorname{Rg}(\mathcal{F}_{1}\otimes\mathcal{F}_{2})$ $=\operatorname{Rg}(\mathcal{F}_{1})\otimes\operatorname{Rg}(\mathcal{F}_{2}),$ (19)
+$\operatorname{Rg}(\mathcal{F}_{1}^{*}\otimes\mathcal{F}_{2}^{*})$ $=\operatorname{Rg}(\mathcal{F}_{1}^{*})\otimes\operatorname{Rg}(\mathcal{F}_{2}^{*}),$ (20)
+$\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}$ $=\widehat{\mathcal{F}_{1}}\otimes\widehat{\mathcal{F}_{2}}$, (21)
+$\lambda(\mathcal{F}_{1}\otimes\mathcal{F}_{2})$ $=\lambda(\mathcal{F}_{1})\otimes\lambda(\mathcal{F}_{2}),$ (22)
+where $\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}$, $\widehat{\mathcal{F}}_{2}$ and $\widehat{\mathcal{F}}_{2}$ are respectively the reduced channels associated to $\mathcal{F}_{1}\otimes\mathcal{F}_{2}$, $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$.
 Proof.
-Let $\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$ and $\widetilde{\mathcal{H}}=\bigoplus_{k=1}^{K}\widetilde{\mathcal{H}}_{k,1}\otimes\widetilde{\mathcal{H}}_{k,2}$ be such that
-[…]
-[…]
-Then, let
-$\mathcal{E}:x\in\mathcal{L}(\mathcal{H})\mapsto\sum_{k=1}^{K}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\tilde{\omega}_{k}\text{ and }\mathcal{D}:x\in\mathcal{L}(\widetilde{\mathcal{H}})\mapsto\sum_{k=1}^{K}U_{k}^{*}\widetilde{\operatorname{\operatorname{tr}}}_{k,2}(\widetilde{P}_{k}x\widetilde{P}_{k})U_{k}\otimes\rho_{k}$
-with a set of states $\{\widetilde{\omega}_{k}\in\mathcal{D}(\widetilde{\mathcal{H}}_{k,2})~|~k\in[K]\}$ and, for all $k\in[K]$, a unitary $U_{k}:\mathcal{H}_{k,1}\mapsto\widetilde{\mathcal{H}}_{k,1}$.
-These unitaries allow to identify $\mathcal{L}(\mathcal{H}_{k,1})$ with $\mathcal{L}(\widetilde{\mathcal{H}}_{k,1})$ for each $k\in[K]$.
-Then, for $x\in\mathcal{L}(\mathcal{H})$, we have for any $k,k^{\prime}\in[K]$
-$\widetilde{\operatorname{\operatorname{tr}}}_{k^{\prime},2}(\widetilde{P}_{k^{\prime}}(U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\tilde{\omega}_{k})\widetilde{P}_{k^{\prime}})=\delta_{k=k^{\prime}}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}.$
-As a result,
-$\widetilde{\mathcal{F}}(\mathcal{E}(x))=\sum_{k=1}^{K}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\widetilde{\rho}_{k}.$
-It follows that
-$\mathcal{D}(\widetilde{\mathcal{F}}(\mathcal{E}(x)))$ $=\sum_{k=1}^{K}U_{k}^{*}\widetilde{\operatorname{\operatorname{tr}}}_{k,2}(U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\widetilde{\rho}_{k})U_{k}\otimes\rho_{k}$
-$=\sum_{k=1}^{K}U_{k}^{*}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}U_{k}\otimes\rho_{k}$
-$=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}=\mathcal{F}(x).$
-We now show that the range and the shape vector of idempotent channels preserve the tensor product, which property will be used in the proofs of our Theorems 1 and 2.
-Proposition 4.
-For $\mathcal{F}$ an idempotent channel, the range of $\mathcal{F}^{*}$ and thus the shape of $\mathcal{F}$ satisfy for all $n\in\mathbb{N}^{*}$:
-$\operatorname{Rg}(\mathcal{F}^{*\otimes n})=\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n},$ (13)
-$\lambda(\mathcal{F}^{\otimes n})=\lambda(\mathcal{F})^{\otimes n}.$ (14)
-The rest of this section is devoted to proving Propositions 1 and 4.
-We begin by proving Proposition 1. We remark that the set of fixed points and the range of an idempotent function coincide.
-Lemma 3.
-Let $f:\Omega\rightarrow\Omega$ be an arbitrary idempotent function, with $\Omega$ an arbitrary set, then
-$\operatorname{Rg}(f)\coloneqq\{f(x)~|~x\in\Omega\}=\{x\in\Omega~|~f(x)=x\}.$ (15)
-Proof.
-If $f(x)=x$, then $f(f(x))=f(x)=x$, so that $x\in\operatorname{Rg}(f)$. For the other inclusion, if $y=f(x)$, we also have $f(y)=f(f(x))=f(x)=y$, so $y$ is a fixed point. ∎
-Proposition 1 basically follows from statements in [18] and we detail it here only for the convenience of the reader.
-Proof of Proposition 1.
-[…]
-The structure of these algebras is well-understood (see for example [16, 5, 6]).
-In finite dimensions, any *-algebra can be decomposed as a direct sum of algebras of complex square matrices (which, in the theory of operator algebras, are called factors [16, 5, 6]).
-This allows us to write
-$\operatorname{Rg}(\mathcal{F}^{*})\cong\bigoplus_{k=1}^{K}\mathcal{L}({\mathbb{C}}^{d_{k}}),$ (16)
-for a set of integers $\{d_{1},\dots,d_{K}\}$, where $\cong$ denotes a *-algebra isomorphism.
-Furthermore, this decomposition is unique up to permutation of the terms in the right-hand side of eq.
-(16) and characterizes uniquely each *-algebra, in the sense that two *-algebras are isomorphic if and only if they have the same decomposition up to permutation of the terms in the direct sum.
-We can go even further and give the general form of $\mathcal{F}^{*}$ using Proposition 1.5 in [18].
-Indeed, we can decompose the Hilbert space $\mathcal{H}$ as a direct sum $\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, where we write $d_{k}=\dim(\mathcal{H}_{k,1})$ and $m_{k}=\dim(\mathcal{H}_{k,2})$, define a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$ and a set of orthogonal projectors $\{P_{k}~|~\operatorname{Rg}(P_{k})=\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},k\in[K]\}$ such that $\mathcal{F}^{*}$ takes the form:
-$\forall x\in\mathcal{L}(\mathcal{H}),~\mathcal{F}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}}.$ (17)
-The range $\operatorname{Rg}(\mathcal{F}^{*})$ will then be decomposed as
-$\operatorname{Rg}(\mathcal{F}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}.$ (18)
-Indeed, let $x=\sum_{k=1}^{K}x_{k}\otimes\mathds{1}_{m_{k}}$ with $x_{k}\in\mathcal{L}(\mathcal{H}_{k,1})$ for all $k\in[K]$, then:
-$\mathcal{F}^{*}(x)$ $=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(x_{k}\otimes\rho_{k})\otimes\mathds{1}_{m_{k}}=x,$
-so that $x\in\operatorname{Rg}(\mathcal{F}^{*})$.
-This proves eq.
-(10).
-Note that eq.
-(18) implies eq.
-(16) as each term $\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}$ is isomorphic, as a *-algebra, to $\mathcal{L}(\mathcal{H}_{k,1})$.
-Then, as $\mathcal{F}$ is the adjoint of $\mathcal{F}^{*}$, we can take the adjoint of eq. (17) and get:
-$\forall x\in\mathcal{L}(\mathcal{H}),~\mathcal{F}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}.$ (19)
-Thus, whenever $\mathcal{F}$ is an idempotent channel, its range admits the decomposition
-$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$ (20)
-which proves eq. (11). ∎
-We then proceed to the proof of Proposition 4, which will allow us to reduce the study of the problem of emulating $k$ copies of $\mathcal{F}$ using $n$ copies of $\mathcal{G}$ for general pairs of integers $(k,n)$ to the case where $k=n$.
-The fact that we can do such a reduction is probably the reason why we get a single-letter formula for the emulation capacity.
-Proof of Proposition 4.
-As, by definition, $\lambda(\mathcal{F})$ is obtained from the decomposition in factors of the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$, it follows that eq. (14) is a consequence of eq. (13).
-Let $n\in\mathbb{N}^{*}$, we take the $n$-th tensor power of eq. (17) to get that, for $x\in\mathcal{L}(\mathcal{H})$,
-$\mathcal{F}^{*\otimes n}(x)=\sum_{\bar{k}\in[K]^{n}}\operatorname{\operatorname{tr}}_{\bar{k},2}(P_{\bar{k}}xP_{\bar{k}}(\mathds{1}_{d_{\bar{k}}}\otimes\rho_{\bar{k}}))\otimes\mathds{1}_{m_{\bar{k}}},$ (21)
-with $\bar{k}=(k_{1},\dots,k_{n})$ ranging on all the $n$-uples in $K]^{n}$, $\operatorname{\operatorname{tr}}_{\bar{k},2}=\operatorname{\operatorname{tr}}_{k_{1},2}\otimes\dots\otimes\operatorname{\operatorname{tr}}_{k_{n},2}$ are the product partial traces, $P_{\bar{k}}=P_{k_{1}}\otimes\dots\otimes P_{k_{n}}$ are the product projections, $\rho_{\bar{k}}=\rho_{k_{1}}\otimes\dots\otimes\rho_{k_{n}}$ are product states, and similarly $\mathds{1}_{d_{\bar{k}}}=\mathds{1}_{d_{k_{1}}}\otimes\dots\otimes\mathds{1}_{d_{k_{n}}}$.
-Hence, the range of $\mathcal{F}^{*}$ can be decomposed as:
-$\operatorname{Rg}(\mathcal{F}^{*\otimes n})=\bigoplus_{\bar{k}\in[K]^{n}}\mathcal{L}(\mathcal{H}_{\bar{k}})\otimes\mathds{1}_{m_{\bar{k}}}=\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n},$ (22)
-which completes the proof of Proposition 4. ∎
+The proof of eq.
+(19) and eq.
+(20) is elementary and these equations actually hold for every linear maps $\Phi:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{1}^{\prime})$ and $\Psi:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{L}(\mathcal{H}_{2}^{\prime})$.
+Indeed, since $\operatorname{Rg}(\Phi)\otimes\operatorname{Rg}(\Psi)=\operatorname{\operatorname{Span}}\{y_{1}\otimes y_{2}~|~y_{1}\in\operatorname{Rg}(\Phi),y_{2}\in\operatorname{Rg}(\Psi)\}$, we argue on product elements $y_{1}\otimes y_{2}\in\operatorname{Rg}(\Phi)\otimes\operatorname{Rg}(\Psi)$ with $y_{1}\in\operatorname{Rg}(\Phi)$ and $y_{2}\in\operatorname{Rg}(\Psi)$.
+We have
+$y_{1}\otimes y_{2}=\Phi(x_{1})\otimes\Psi(x_{2})=\Phi\otimes\Psi(x_{1}\otimes x_{2})\in\operatorname{Rg}(\Phi\otimes\Psi),$
+where $x_{1}\in\mathcal{L}(\mathcal{H}_{1})$ and $x_{2}\in\mathcal{L}(\mathcal{H}_{2})$.
+Conversely, for $y\in\operatorname{Rg}(\Phi\otimes\Psi)$ we have $y=\Phi\otimes\Psi(x)$, with $x\in\mathcal{L}(\mathcal{H}_{1}\otimes\mathcal{H}_{2})$.
+Thus, we can write $x$ as
+$x=\sum_{j=1}^{J}x_{1}^{(j)}\otimes x_{2}^{(j)},$
+with $x_{i}^{(j)}\in\mathcal{L}(\mathcal{H}_{i})$ for $(i,j)\in[2]\times[J]$. Therefore,
+$y=\Phi\otimes\Psi(\sum_{j=1}^{J}x_{1}^{(j)}\otimes x_{2}^{(j)})=\sum_{j=1}^{J}\Phi(x_{1}^{(j)})\otimes\Psi(x_{2}^{(j)})\in\operatorname{Rg}(\Phi)\otimes\operatorname{Rg}(\Psi).$
+Then, we prove the compatibility of taking the reduced version of idempotent channels with the tensor product.
+We write $\mathds{1}_{1}$ and $\mathds{1}_{2}$ respectively the identity operators in $\mathcal{L}(\mathcal{H}_{1})$ and $\mathcal{L}(\mathcal{H}_{2})$ and $\mathds{1}_{1,2}=\mathds{1}_{1}\otimes\mathds{1}_{2}$.
+We also write $\mathcal{H}_{i}^{(0)}=\operatorname{\operatorname{supp}}(\mathcal{F}_{i}(\mathds{1}_{i}))$ and $\mathcal{H}_{i}^{(1)}=\ker(\mathcal{F}_{i}(\mathds{1}_{i}))$ for $i\in[2]$ and the corresponding isometries $V_{i}:\mathcal{H}_{i}^{(0)}\rightarrow\mathcal{H}_{i}$ defining $\widehat{\mathcal{F}}_{i}$.
+Then, $\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2})=\mathcal{F}_{1}(\mathds{1}_{1})\otimes\mathcal{F}_{2}(\mathds{1}_{2})$, hence $\operatorname{\operatorname{supp}}(\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2}))=\operatorname{\operatorname{supp}}(\mathcal{F}_{1}(\mathds{1}_{1}))\otimes\operatorname{\operatorname{supp}}(\mathcal{F}_{2}(\mathds{1}_{2}))$ and $\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}$ maps $\mathcal{L}(\mathcal{H}_{1}^{(0)}\otimes\mathcal{H}_{2}^{(0)})$ into itself.
+If we write $e_{1,2}=\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2})^{0}$ the orthogonal projector onto $\mathcal{H}_{1}^{(0)}\otimes\mathcal{H}_{2}^{(0)}$ the support of $\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2})$, we also have that $e_{1,2}=e_{1}\otimes e_{2}=(V_{1}\otimes V_{2})(V_{1}\otimes V_{2})^{*}$ with $e_{1}=V_{1}V_{1}^{*}=\mathcal{F}_{1}(\mathds{1}_{1})^{0}$ and $e_{2}=V_{2}V_{2}^{*}=\mathcal{F}_{2}(\mathds{1}_{2})^{0}$ the orthogonal projectors onto the respective supports of $\mathcal{F}_{i}(\mathds{1}_{i})$, $i\in[2]$.
+Then for $x_{1}\in\mathcal{L}(\mathcal{H}_{1})$ and $x_{2}\in\mathcal{L}(\mathcal{H}_{2})$, we have
+$\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}(x_{1}\otimes x_{2})$ $=(V_{1}\otimes V_{2})^{*}\mathcal{F}_{1}\otimes\mathcal{F}_{2}((V_{1}\otimes V_{2})(x_{1}\otimes x_{2})(V_{1}\otimes V_{2})^{*})(V_{1}\otimes V_{2})$
+$=V_{1}^{*}\mathcal{F}_{1}(V_{1}x_{1}V_{1}^{*})V_{1}\otimes V_{2}^{*}\mathcal{F}_{2}(V_{2}x_{2}V_{2}^{*})V_{2}$
+$=\widehat{\mathcal{F}}_{1}(x_{1})\otimes\widehat{\mathcal{F}}_{2}(x_{2}).$
+Eq. (21) then follows on the whole space $\mathcal{L}(\mathcal{H}_{1}\otimes\mathcal{H}_{2})=\mathcal{L}(\mathcal{H}_{1})\otimes\mathcal{L}(\mathcal{H}_{2})$ by linearity.
+The last equality, eq. (22), is then seen by injecting eq. (8) into eq. (19) and by the definition of the shape vector of an idempotent channel. ∎
 We end this section by the following remark on the origin of the concept of shape vectors.
 Remark 1.
-Note that the shape of an idempotent channel $\mathcal{F}$ coincides with the shape of the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$.
+Note that the shape of an idempotent channel $\mathcal{F}$ coincides with the shape of the *-algebra $\operatorname{Rg}(\widehat{\mathcal{F}}^{*})$, with $\widehat{\mathcal{F}}$ the reduced channel associated to $\mathcal{F}$.
 The notion of shape vector for *-algebras was introduced by Kuperberg in [12].
@@
 for a set of integers $\{d_{k}~|~k\in[K]\}$ and we write $\lambda(\mathcal{A})$ the shape vector of $\mathcal{A}$, which coordinates are the integers $(d_{k})_{k\in[K]}$ sorted in non-increasing order.
-[^2] In fact, in all this article and in Kuperberg’s [11], the order with respect to which the coordinates of shape vectors are sorted is actually irrelevant.
-4 Emulation of idempotent channel with zero error
+[^3] In fact, in all this article and in Kuperberg’s work [12], the order with respect to which the coordinates of shape vectors are sorted is actually irrelevant.
+4 Emulation of idempotent channels with zero error
 Now that we have presented the decomposition properties of idempotent channels, we move on to proving Theorem 1 in this section, which is divided into two subsections.
@@
 $\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}$ (23)
-is achievable. In the second one, we prove that there is no rate strictly greater than eq. (23).
+is achievable. In the second one, we prove that any rate strictly greater than eq. (23) is not achievable.
 Before presenting these proofs, we provide some examples for the expression eq.
@@
 Note that in the first case, i.e.
-$\lambda(\mathcal{F})$ has a coefficient greater or equal to two, the emulation capacity of $\mathcal{F}$ by the completely dephasing channel is zero, that is that it is impossible to emulate a channel preserving some quantum information with a classical one, without error.
+$\lambda(\mathcal{F})$ has a coefficient greater or equal to two, the emulation capacity of $\mathcal{F}$ by the completely dephasing channel is zero, that is it is impossible to emulate a channel preserving some quantum information with a classical one, without error.
 $\mathcal{F}$ $\mathcal{G}$ $C(\mathcal{G}\mapsto\mathcal{F})$
@@
 $C(\mathcal{G}\mapsto\mathcal{F})^{-1}=2\neq 1=C(\mathcal{F}\mapsto\mathcal{G}).$
-This example shows that neither the emulation capacity nor the strong converse rate given in Theorem 2 are reversible, since, in this example, the emulation capacity matches with the strong converse rate of Theorem 2, as $\mathcal{F}$ is taken to be an identity channel.
+This example shows that neither the emulation capacity nor the strong converse rate given in Theorem 2 are reversible, since, as $\mathcal{F}$ is taken to be an identity channel in this example, both quantity match.
 Although, when either $\mathcal{F}$ or $\mathcal{G}$ is an identity channel or a completely dephasing one, the infimum in eq.
@@
 Proof.
-This theorem, despite not stated explicitly in [11], is proven in the proof of Theorem 1.1 of [11]. ∎
+This theorem, despite not being stated explicitly in [12], is proven in the proof of Theorem 1.1 of [12]. ∎
 For our achievability proof, it will be useful to lift $\iota$ into a unital completely positive map.
+In the statement of the following lemma, we will use the fact that it is always possible, given $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ with the same identity operator as $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$, to construct a unital completely positive map $E_{\mathcal{A}}:\mathcal{L}(\mathcal{H}_{\mathcal{A}})\rightarrow\mathcal{A}$ such that for all $x\in\mathcal{A}$, $E_{\mathcal{A}}(x)=x$.
+This fact is easily seen as if $\mathcal{A}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ with the same identity operator, we can write the orthogonal direct sum decompositions
+$\mathcal{H}_{\mathcal{A}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},\quad\mathcal{A}=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}},$
+with $m_{k}=\dim(\mathcal{H}_{k,2})$ for $k\in[K]$. Then, we can take $E_{\mathcal{A}}:\mathcal{L}(\mathcal{H}_{\mathcal{A}})\rightarrow\mathcal{A}$ defined by
+$E_{\mathcal{A}}:x\rightarrow\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\frac{1}{m_{k}}\mathds{1}_{m_{k}},$ (26)
+which is called the tracial conditional expectation on $\mathcal{A}$ (see e.g. Proposition 1.5 in [19]).
 Lemma 6.
-Let $\mathcal{A}$, $\mathcal{B}$ be two finite-dimensional *-algebras and $\iota:\mathcal{A}\rightarrow\mathcal{B}$ be a subunital *-homomorphism.
-Let $\mathds{1}_{\mathcal{A}}\in\mathcal{A}$ and $\mathds{1}_{\mathcal{B}}\in\mathcal{B}$ be the identity operator of $\mathcal{A}$ and $\mathcal{B}$ respectively.
-Then, for all $x\in\mathcal{A}$,
-$\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}})).$ (26)
-Furthermore, $P_{\iota}\coloneqq\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector, and the map
-$\widetilde{\iota}:x\mapsto\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})$ (27)
-is unital completely positive. Finally, under the assumption that $\iota$ is injective, we write $\iota^{-1}:\iota(\mathcal{A})\subseteq\mathcal{B}\rightarrow\mathcal{A}$ its partial inverse, then
-$\widetilde{\iota}^{-1}:x\in\mathcal{B}\mapsto\iota^{-1}(P_{\iota}xP_{\iota})$ (28)
-is a unital completely positive map such that
-$\forall x\in\mathcal{A},~\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)=x.$ (29)
-Note that none of the two maps $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are *-homomorphisms unless $\iota$ is unital.
+Let $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{A}})$, $\mathcal{B}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ be two finite-dimensional unital *-subalgebras of $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ respectively, where we assume that the identity operator in $\mathcal{A}$ coincides with the one in $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$, denoted $\mathds{1}_{\mathcal{A}}$ and the one in $\mathcal{B}$ coincides with the one in $\mathcal{L}(\mathcal{H}_{\mathcal{B}})$, denoted $\mathds{1}_{\mathcal{B}}$.
+Let $\iota:\mathcal{A}\rightarrow\mathcal{B}$ be an injective subunital *-homomorphism.
+Then, writing $\mathcal{H}_{\mathcal{B}}^{(0)}=\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))\subseteq\mathcal{H}_{\mathcal{B}}$, we have that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})\in\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ is the orthogonal projector on $\mathcal{H}_{\mathcal{B}}^{(0)}$.
+We denote $V:\mathcal{H}_{\mathcal{B}}^{(0)}\rightarrow\mathcal{H}_{\mathcal{B}}$ the isometry embedding $\mathcal{H}_{\mathcal{B}}^{(0)}$ into $\mathcal{H}_{\mathcal{B}}$ such that $VV^{*}=P_{\iota}$ and we define the reduced *-homomorphism $\widehat{\iota}:\mathcal{A}\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ by
+$\widehat{\iota}:x\mapsto V^{*}\iota(x)V.$ (27)
+Then $\widehat{\iota}:\mathcal{A}\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ is a unital injective *-homomorphism and $\widehat{\iota}(\mathcal{A})$ is a unital *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ where both $\widehat{\iota}(\mathcal{A})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ share the same identity operator.
+We will write $\widehat{\iota}^{-1}:\widehat{\iota}(\mathcal{A})\rightarrow\mathcal{A}$ the inverse of $\widehat{\iota}$, which is itself a unital injective *-homomorphism.
+We write $E_{\mathcal{A}}:\mathcal{L}(\mathcal{H}_{\mathcal{A}})\rightarrow\mathcal{A}$ and $E_{\widehat{\iota}(\mathcal{A})}:\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})\rightarrow\widehat{\iota}(\mathcal{A})$ the tracial conditional expectations on $\mathcal{A}$ and $\widehat{\iota}(\mathcal{A})$ respectively.
+Then, the maps
+[…]
+$x$ $\mapsto\iota(E_{\mathcal{A}}(x))+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota}),$
+and
+$\widetilde{\iota}^{-1}:\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ $\rightarrow\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ (29)
+$x$ $\mapsto\widehat{\iota}^{-1}(E_{\widehat{\iota}(\mathcal{A})}(V^{*}xV))$
+are unital completely positive and satisfy for all $x\in\mathcal{A}$,
+$\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)=x.$ (30)
 Proof.
-We start by proving eq. (26). Let $x\in\mathcal{A}$, we can use the polarisation identity (eq. 1.45 in [18]) to write
-$x=\frac{1}{4}\sum_{k=0}^{3}i^{k}(x+i^{k}\mathds{1}_{\mathcal{A}})^{*}(x+i^{k}\mathds{1}_{\mathcal{A}}),$
-so that $x$ is decomposed as a linear combination of four positive operators in $\mathcal{A}$.
-Thus, it suffices to show eq. (26) for positive semidefinite operators. For positive semidefinite $x\in\mathcal{A}$,
-$0\leq x\leq\|x\|\mathds{1}_{\mathcal{A}}.$
-Since $\iota$ is completely positive,
-$0\leq\iota(x)\leq\|x\|\iota(\mathds{1}_{\mathcal{A}}).$
-Thus, $\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$. Now,
-$\iota(\mathds{1}_{\mathcal{A}})^{2}=\iota(\mathds{1}_{\mathcal{A}}^{2})=\iota(\mathds{1}_{\mathcal{A}}),$
-and, as $\iota$ is a *-homomorphism, it preserves the *-operation on $\mathcal{A}$, so that
-$\iota(\mathds{1}_{\mathcal{A}})^{*}=\iota(\mathds{1}_{\mathcal{A}}).$
-This shows that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector in $\mathcal{B}$.
-For $\widetilde{\iota}$, note that, as $\iota$ is subunital, $P_{\iota}\leq\mathds{1}_{\mathcal{B}}$, so that $\mathds{1}_{\mathcal{B}}-P_{\iota}\geq 0$ and $x\mapsto\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})$ is completely positive.
-The unitality is trivial.
-Then, as $\iota^{-1}$ is a *-homomorphism, it is completely positive.
-Furthermore, as $P_{\iota}$ is an orthogonal projector, $x\in\mathcal{B}\mapsto P_{\iota}xP_{\iota}$ is completely positive.
-Therefore the composition $x\mapsto\iota^{-1}(P_{\iota}xP_{\iota})$ is completely positive as well.
-Then $\widetilde{\iota}^{-1}(\mathds{1}_{\mathcal{B}})=\iota^{-1}(P_{\iota})=\iota^{-1}(\iota(\mathds{1}_{\mathcal{A}}))=\mathds{1}_{\mathcal{A}}$.
-Finally, for $x\in\mathcal{A}$, we have
-$\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)$ $=\iota^{-1}(P_{\iota}\bigl(\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})\bigr)P_{\iota})$
-$=\iota^{-1}(P_{\iota}\iota(x)P_{\iota})$
-$=x,$
-where we used for the last equality that $\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(P_{\iota})$ so that $P_{\iota}\iota(x)P_{\iota}=\iota(x)$. ∎
+First, we prove that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector on $\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$.
+As $\iota$ is a *-homomorphism, it preserves both the operation of taking the adjoint and the multiplication, so that $P_{\iota}^{*}=\iota(\mathds{1}_{\mathcal{A}})^{*}=\iota(\mathds{1}_{\mathcal{A}}^{*})=P_{\iota}$ and $P_{\iota}^{2}=\iota(\mathds{1}_{\mathcal{A}})^{2}=\iota(\mathds{1}_{\mathcal{A}}^{2})=P_{\iota}$.
+We move on to prove that $\widehat{\iota}:\mathcal{A}\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$, with $\mathcal{H}_{\mathcal{B}}^{(0)}=\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$ is a *-homomorphism, which is also unital and injective.
+First $\widehat{\iota}$ is clearly linear.
+Let $x\in\mathcal{A}$,
+$\widehat{\iota}(x)^{*}=(V^{*}\iota(x)V)^{*}=V\iota(x)^{*}V^{*}=V\iota(x^{*})V^{*}=\widehat{\iota}(x^{*}),$
+and for $y\in\mathcal{A}$,
+$\widehat{\iota}(xy)=V^{*}\iota(xy)V=V^{*}\iota(x)\iota(y)V\overset{(a)}{=}V^{*}VV^{*}\iota(x)VV^{*}\iota(y)V=\widehat{\iota}(x)\widehat{\iota}(y).$
+The equality $(a)$ holds by applying Lemma 3 on the map $\Phi=\iota$ which is positive as it is a *-homomorphism [^4] Actually, *-homomorphisms are even completely positive [7]..
+To show the unitality, we have
+$\widehat{\iota}(\mathds{1}_{\mathcal{A}})=V^{*}\iota(\mathds{1}_{\mathcal{A}})V=V^{*}P_{\iota}V=V^{*}VV^{*}V=\mathds{1}_{\mathcal{H}_{\mathcal{B}}^{(0)}},$
+[…]
+$0=\widehat{\iota}(x)=V^{*}\iota(x)V.$
+Hence, by Lemma 3
+$\iota(x)=VV^{*}\iota(x)VV^{*}=0,$
+so that $x\in\ker(\iota)$, thus $x=0$ by injectivity of $\iota$.
+Since $\widehat{\iota}$ is a unital *-homomorphism, $\widehat{\iota}(\mathcal{A})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ and has the same identity operator as $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$, so that the tracial conditional expectation $E_{\widehat{\iota}(\mathcal{A})}$ is well-defined by eq.
+(26).
+Furthermore, as $\widehat{\iota}$ is injective $\widehat{\iota}^{-1}$ is well-defined and is a unital *-homomorphism as well.
+Now, we show that $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are unital completely positive maps.
+By unitality of $E_{\mathcal{A}}$, $E_{\widehat{\iota}(\mathcal{A})}$ and $\widehat{\iota}^{-1}$, it is clear that $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are unital.
+Furthermore, $\widetilde{\iota}^{-1}$ is the composition of completely positive maps and is thus completely positive.
+Moreover, as $\iota$ was assumed to be subunital, $\iota(\mathds{1}_{\mathcal{A}})=P_{\iota}\leq\mathds{1}_{\mathcal{B}}$ so that $\widetilde{\iota}$ is completely positive.
+We also have that $\widetilde{\iota}(\mathcal{L}(\mathcal{H}_{\mathcal{A}}))\subseteq\mathcal{B}$.
+Indeed, as $\mathcal{B}$ is unital of unit $\mathds{1}_{\mathcal{B}}$ and as the range of $\iota$ is in $\mathcal{B}$, $\mathds{1}_{\mathcal{B}}-P_{\iota}\in\mathcal{B}$ and $\iota(E_{\mathcal{A}}(x))\in\mathcal{B}$ for all $x\in\mathcal{L}(\mathcal{H}_{\mathcal{B}})$.
+Hence $\widetilde{\iota}(\mathcal{A})\subseteq\mathcal{B}$.
+Finally, note that $\mathds{1}_{\mathcal{B}}-P_{\iota}$ is the orthogonal projector on $\ker(P_{\iota})$ which is orthogonal to $\operatorname{\operatorname{supp}}(P_{\iota})=\operatorname{Rg}(V)$, so that $(\mathds{1}_{\mathcal{B}}-P_{\iota})V=0$.
+Hence, for $x\in\mathcal{A}$,
+$\widetilde{\iota}^{-1}(\widetilde{\iota}(x))=\widehat{\iota}^{-1}(E_{\widehat{\iota}(\mathcal{A})}(V^{*}\iota(E_{\mathcal{A}}(x))V))=\widehat{\iota}^{-1}(E_{\widehat{\iota}(\mathcal{A})}(\widehat{\iota}(x)))=\widehat{\iota}^{-1}(\widehat{\iota}(x))=x.$
 We now prove Proposition 6.
@@
 $\mathcal{F}^{*\otimes n}=\widetilde{\iota}^{-1}\circ\mathcal{G}^{*\otimes n}\circ\widetilde{\iota}\circ\mathcal{F}^{*\otimes n},$ (31)
-with $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ being defined as in Lemma 5. Indeed, for $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$,
-$\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\iota(\mathcal{F}^{*\otimes n}(x))+\frac{\operatorname{\operatorname{tr}}(\mathcal{F}^{*\otimes n}(x))}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{F}^{\otimes n}})}(\mathds{1}_{\mathcal{G}^{\otimes n}}-\iota(\mathds{1}_{\mathcal{F}^{\otimes n}}))\in\operatorname{Rg}(\mathcal{G}^{*\otimes n}),$
-where $\mathds{1}_{\mathcal{F}^{\otimes n}}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$ and $\mathds{1}_{\mathcal{G}^{\otimes n}}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$.
-Since $\mathcal{G}^{*\otimes n}$ is idempotent, every element of its range is one of its fixed points, so that $\mathcal{G}^{*\otimes n}(\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x)))=\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))$.
-We finally have by eq.
-(29) that $\widetilde{\iota}^{-1}\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\mathcal{F}^{*\otimes n}(x)$.
+with $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ being defined as in Lemma 6.
+Indeed, for $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$, $\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))\in\operatorname{Rg}(\mathcal{G}^{*\otimes n})$.
+Then, since $\mathcal{G}^{*\otimes n}$ is idempotent, every element of its range is one of its fixed points, so that $\mathcal{G}^{*\otimes n}(\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x)))=\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))$.
+As $\mathcal{F}^{*\otimes n}(x)\in\operatorname{Rg}(\mathcal{F}^{*\otimes n})$, we can finally apply eq.
+(30) to get that $\widetilde{\iota}^{-1}\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\mathcal{F}^{*\otimes n}(x)$.
 We then write $\mathcal{E}^{*}=\widetilde{\iota}^{-1}$ and $\mathcal{D}^{*}=\widetilde{\iota}\circ\mathcal{F}^{*\otimes n}$.
@@
 $\|\lambda(\mathcal{A})\|_{p}\leq\|\lambda(\mathcal{B})\|_{p}.$ (32)
-Note that Theorems 4 and 6 are not converse one from another because the inequality between the $\ell_{p}$-norms of the shape vectors is strict in one case but not in the other.
+Note that Theorems 5 and 7 are not converse one from another because the inequality between the $\ell_{p}$-norms of the shape vectors is strict in one case but not in the other.
 The converse of Theorem 1 then follows from Theorem 7 combined with the following proposition.
@@
 $\mathcal{F}=\mathcal{D}\mathcal{G}\mathcal{E}.$ (33)
-We write $\mathds{1}_{\mathcal{F}}$ and $\mathds{1}_{\mathcal{G}}$ the identity operator in $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$, respectively.
+We write $\mathds{1}_{\mathcal{F}}$ and $\mathds{1}_{\mathcal{G}}$ the identity operator in $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ respectively.
 To simplify our notations, we also denote $e_{\mathcal{G}\mathcal{E}}=\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})^{0}$.
@@
 $\widetilde{\mathcal{G}}^{*}:x\mapsto e_{\mathcal{G}\mathcal{E}}\mathcal{G}^{*}(x)e_{\mathcal{G}\mathcal{E}}.$ (34)
-Then, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism which embeds $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ as a subalgebra.
+Then, $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a unital *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ whose identity operator is $e_{\mathcal{G}\mathcal{E}}$ and $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism which embeds $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ as a subalgebra.
 Furthermore, for all $p\in[1,+\infty]$
@@
 Lemma 8.
-Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be a quantum channel and write $e=\Phi(\mathds{1})^{0}$. Then
+Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be a quantum channel and write $e=\Phi(\mathds{1})^{0}$. Then,
   1. i)
-    For all $x\in\mathcal{L}(\mathcal{H}^{\prime})$, $\Phi^{*}(x)=\Phi^{*}(exe)$.
+    for all $x\in\mathcal{L}(\mathcal{H}^{\prime})$, $\Phi^{*}(x)=\Phi^{*}(exe)$,
   2. ii)
@@
 Proof.
-Let $y\in\mathcal{L}(\mathcal{H})$ be positive semidefinite. By positivity of $\Phi$,
-$\Phi(y)\leq\|y\|\Phi(\mathds{1}).$
-Therefore, $\operatorname{\operatorname{supp}}(\Phi(y))\subseteq\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))=e$. Then, for all $x\in\mathcal{L}(\mathcal{H}^{\prime})$,
-$\operatorname{\operatorname{tr}}(y\Phi^{*}(x))=\operatorname{\operatorname{tr}}(\Phi(y)x)=\operatorname{\operatorname{tr}}(e\Phi(y)ex)=\operatorname{\operatorname{tr}}(y\Phi^{*}(exe)).$
-By linearity, this equation extends to all $y\in\mathcal{L}(\mathcal{H})$. Thus, $\Phi^{*}(x)=\Phi^{*}(exe)$.
+Let $\mathcal{H}_{0}^{\prime}=\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))$ and $V:\mathcal{H}_{0}^{\prime}\rightarrow\mathcal{H}^{\prime}$ be the isometry such that $e=VV^{*}$.
+Let $y\in\mathcal{L}(\mathcal{H})$ and $x\in\mathcal{L}(\mathcal{H}^{\prime})$, then
+$\operatorname{\operatorname{tr}}(y\Phi^{*}(x))=\operatorname{\operatorname{tr}}(\Phi(y)x)=\operatorname{\operatorname{tr}}(VV^{*}\Phi(y)VV^{*}x)=\operatorname{\operatorname{tr}}(e\Phi(y)ex)=\operatorname{\operatorname{tr}}(y\Phi^{*}(exe)),$
+where the second equality follows by Lemma 3. Therefore $\Phi^{*}(x)=\Phi(exe)$.
 For the second statement, let $x\in\mathcal{L}(\mathcal{H}^{\prime})$ be positive semidefinite such that $\Phi^{*}(x)=0$. Then
@@
 $\{x\in\mathcal{L}(\mathcal{H}^{\prime})~|~\mathcal{F}^{*}(x^{*}x)=\mathcal{F}^{*}(x^{*})\mathcal{F}^{*}(x)\text{ and }\mathcal{F}^{*}(xx^{*})=\mathcal{F}^{*}(x)\mathcal{F}^{*}(x^{*})\}.$
-In particular, if $\mathcal{F}^{*}$ is subunital completely positive [^3] In fact, it suffices for $\mathcal{F}^{*}$ to be subunital $2$-positive for it to satisfy the Kadison-Schwarz inequality, see Section 5.2 of [18] for more details., it satisfies the Kadison-Schwarz inequality (eq.
+In particular, if $\mathcal{F}^{*}$ is unital completely positive [^5] In fact, it suffices for $\mathcal{F}^{*}$ to be subunital $2$-positive for it to satisfy the Kadison-Schwarz inequality, see Section 5.2 of [19] for more details., it satisfies the Kadison-Schwarz inequality (eq.
 (35)) and we can use this second characterisation of its multiplicative domain.
@@
 Let $x\in\operatorname{Rg}(\mathcal{F}^{*})$.
-As $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-algebra, $x^{*}x\in\operatorname{Rg}(\mathcal{F}^{*})$, and as $\mathcal{F}^{*}$ is idempotent, by Lemma 3, we have:
+By the assumption that $\mathcal{F}=\widehat{\mathcal{F}}$, $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$, thus $x^{*}x\in\operatorname{Rg}(\mathcal{F}^{*})$, and as $\mathcal{F}^{*}$ is idempotent, $xx^{*}$ is a fixed point of $\mathcal{F}^{*}$, so that we have:
 $x^{*}x$ $=\mathcal{F}^{*}(x^{*}x)$
@@
 Therefore, $x$ is in the multiplicative domain of $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$.
-Thus, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism into the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$.
+Thus, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism into the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$, which is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ and whose identity operator is $e_{\mathcal{G}\mathcal{E}}$.
 It is injective as, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
-$x=\mathcal{F}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x).$
+$x=\mathcal{F}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x),$
+hence $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)=0$ implies $x=0$.
 It is unital as it maps the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ to $e_{\mathcal{G}\mathcal{E}}$ which is the identity operator of the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$.
 Indeed, as both $\mathcal{G}$ and $\mathcal{D}$ are channels, $\mathcal{G}^{*}$ and $\mathcal{D}^{*}$ are unital so that $\mathcal{G}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=\mathds{1}_{\mathcal{G}}$ and $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=e_{\mathcal{G}\mathcal{E}}$.
-Now, to show that for all $p\in[1,+\infty]$, $\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}$, it suffices to remark that, by Proposition 1, we can decompose the Hilbert space $\mathcal{H}_{\mathcal{G}}$ as
+Now, to show that $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ which satisfies for all $p\in[1,+\infty]$, $\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}$}, it suffices to remark that, by Proposition 1, we can decompose the Hilbert space $\mathcal{H}_{\mathcal{G}}$ as
 $\mathcal{H}_{\mathcal{G}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$
@@
 $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}=\bigoplus_{k=1}^{K}x_{k}^{0}\mathcal{L}(\mathcal{H}_{k,1})x_{k}^{0}\otimes\rho_{k}^{0}.$
-Hence, for all $k\in[K]$, the $k$-th coordinate of the shape vector of $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ satisfies
+Hence, $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a *-subalgebra for all $k\in[K]$, the $k$-th coordinate of the shape vector of $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ satisfies
 $\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})_{k}\leq\lambda(\operatorname{Rg}(\mathcal{G}^{*}))_{k},$
@@
 $\mathcal{F}^{\otimes k}=\mathcal{D}\mathcal{G}^{\otimes n}\mathcal{E}.$
-Then by Proposition 6, there is an injective unital *-homomorphism from $\operatorname{Rg}(\mathcal{F}^{*\otimes k})$ into the *-algebra $e_{\mathcal{G}^{\otimes n}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*\otimes n})e_{\mathcal{G}^{\otimes n}\mathcal{E}}$.
+By Proposition 3, then, if we denote $\widehat{\mathcal{F}^{\otimes k}}$ and $\widehat{\mathcal{G}^{\otimes n}}$ the reduced channels associated to $\mathcal{F}^{\otimes k}$ and $\mathcal{G}^{\otimes n}$, this condition is equivalent to
+$\widehat{\mathcal{F}^{\otimes n}}=\mathcal{D}^{\prime}\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime},$
+for some channels $\mathcal{D}^{\prime}$, $\mathcal{E}^{\prime}$.
+Thus, by Proposition 7, there is an injective unital *-homomorphism from $\operatorname{Rg}(\widehat{\mathcal{F}^{\otimes k}}^{*})$ into the *-algebra $e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}}\operatorname{Rg}(\widehat{\mathcal{G}^{\otimes n}}^{*})e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}}$, which is itself a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$.
 Then, by Theorem 7 and the second statement of Proposition 7, we have for all $p\in[1,+\infty]$:
-$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}=\|\lambda(\operatorname{Rg}(\mathcal{F}^{*\otimes k}))\|_{p}\leq\|\lambda(e_{\mathcal{G}^{\otimes n}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*\otimes n})e_{\mathcal{G}^{\otimes n}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*\otimes n}))\|_{p}.$
+$\|\lambda(\widehat{\mathcal{F}^{\otimes k}})\|_{p}=\|\lambda(\operatorname{Rg}(\widehat{\mathcal{F}^{\otimes k}}^{*}))\|_{p}\leq\|\lambda(e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}}\operatorname{Rg}(\widehat{\mathcal{G}^{\otimes n}}^{*})e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\widehat{\mathcal{G}^{\otimes n}}^{*}))\|_{p}=\|\lambda(\widehat{\mathcal{G}^{\otimes n}}^{*})\|.$
 By the multiplicativity property of the shape vectors and the $\ell_{p}$-norms with the tensor product (Proposition 5), we get for all $p\in[1,+\infty]$:
-$\|\lambda(\mathcal{F})\|_{p}^{k}\leq\|\lambda(\mathcal{G})\|^{n}_{p}.$
+$\|\lambda(\mathcal{F})\|_{p}^{k}=\|\lambda(\widehat{\mathcal{F}})\|_{p}^{k}\leq\|\lambda(\widehat{\mathcal{G}})\|_{p}^{k}=\|\lambda(\mathcal{G})\|^{n}_{p}.$
 Thus,
@@
 This stronger result is an application of approximate $C^{*}$-algebras, which were recently introduced by Kitaev in [10].
-5 Channel emulation with errors
+5 Emulation of idempotent channels with errors
 In Theorem 1, the converse states that it is impossible to emulate an idempotent channel $\mathcal{F}$ by another idempotent channel $\mathcal{G}$ at a rate higher than the infimum in eq.
@@
 For any choice of auxiliary system $\mathcal{L}(\mathcal{H}^{\prime\prime})$ of dimension $d=\dim(\mathcal{H}^{\prime\prime})$, for any choice of positive semidefinite operator $\mu\in\mathcal{L}(\mathcal{H}^{\prime}\otimes\mathcal{H}^{\prime\prime})$ satisfying $\mu\leq\mathds{1}$ and of density operator $\sigma\in\mathcal{D}(\mathcal{H}\otimes\mathcal{H}^{\prime\prime})$,
-$\operatorname{\operatorname{tr}}(\mu\Phi_{1}\otimes\mathrm{Id}_{d}(\sigma))-\operatorname{\operatorname{tr}}(\mu\Phi_{2}\otimes\mathrm{Id}_{d}(\sigma))\leq\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}.$ (36)
+$\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}\geq\operatorname{\operatorname{tr}}(\mu\Phi_{1}\otimes\mathrm{Id}_{d}(\sigma))-\operatorname{\operatorname{tr}}(\mu\Phi_{2}\otimes\mathrm{Id}_{d}(\sigma)).$ (37)
 We can now prove Theorem 9.
@@
 […]
+$\inf_{\mathcal{D},\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\inf_{\mathcal{D},\mathcal{E}}\|\widehat{\mathcal{F}}-\mathcal{D}\widehat{\mathcal{G}}\mathcal{E}\|_{\diamond}$.
+Therefore, to simplify our notations, we will suppose without loss of generality that $\mathcal{F}$ and $\mathcal{G}$ are equal to their reduced version, i.e.
+$\mathcal{F}=\widehat{\mathcal{F}}$ and $\mathcal{G}=\widehat{\mathcal{G}}$.
+We prove separately the two following equations:
 $\frac{1}{2}\inf_{\mathcal{D},\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}\geq 1-\frac{\|\lambda(\mathcal{G})\|_{1}}{\|\lambda(\mathcal{F})\|_{1}},$ (38)
@@
 with $p=1$ for one Ansatz and $p=+\infty$ for the other.
-We begin with the case $p=1$. As we saw in Proposition 1, we can decompose $\mathcal{H}_{\mathcal{F}}$ as a direct sum
-$\mathcal{H}_{\mathcal{F}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$
+We begin with the case $p=1$. As we saw in Proposition 1, we can decompose $\mathcal{H}_{\mathcal{F}}$ as an orthogonal direct sum
+$\mathcal{H}_{\mathcal{F}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$
 so that
@@
 $\operatorname{\operatorname{tr}}(\mu\mathcal{D}\mathcal{G}\mathcal{E}\otimes\mathrm{Id}(\sigma))=\frac{1}{\|\lambda(\mathcal{F})\|_{1}}\operatorname{\operatorname{tr}}(\sum_{k=1}^{K}\sum_{\nu=1}^{d_{k}}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})(\mathcal{D}\mathcal{G}\mathcal{E}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))).$ (42)
-From Propososition 1, we can decompose $\mathcal{H}_{\mathcal{G}}$ and $\mathcal{G}$ as
+From Proposition 1, we can decompose $\mathcal{H}_{\mathcal{G}}$ and $\mathcal{G}$ as
 $\mathcal{H}_{\mathcal{G}}=\bigoplus_{k=1}^{K^{\prime}}\mathcal{H}^{\prime}_{k,1}\otimes\mathcal{H}^{\prime}_{k,2},$
@@
 where $(a)$ follows from the fact that $\mathcal{D}\mathcal{G}_{2}:\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ is a channel, so that $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is unital.
-Then $(b)$ follows from the fact that for all $k\in[K]$ and $\nu\in[d_{k}]$, $|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}$ is a state so that $\operatorname{\operatorname{tr}}(\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))=1$, hence $\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}})\leq\mathds{1}_{\mathcal{G}_{1}}$, and that as $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is completely positive, $\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})\geq 0$.
-We then simply use that if $A,B\geq 0$ with $B\leq\mathds{1}$, $\operatorname{\operatorname{tr}}(A)\geq\operatorname{\operatorname{tr}}(AB)$.
+Then $(b)$ follows from the fact that for all $k\in[K]$ and $\nu\in[d_{k}]$, $|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}$ is a state so that $\operatorname{\operatorname{tr}}(\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))=1$, hence $0\leq\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}})\leq\mathds{1}_{\mathcal{G}_{1}}$, and that as $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is completely positive, $\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})\geq 0$.
+We then simply use that if $A\geq 0$ and $B\leq\mathds{1}$, $\operatorname{\operatorname{tr}}(A)\geq\operatorname{\operatorname{tr}}(AB)$.
 When injecting the last equation into eq.
@@
 the case $p=+\infty$.
-To deal with the case where the blocks of the range of $\mathcal{F}^{*}$ would have multiplicities, we factorise $\mathcal{F}=\mathcal{F}_{2}\circ\mathcal{F}_{1}$ as we did for $\mathcal{G}$ in eq.
+To deal with the case where the blocks of the range of $\mathcal{F}^{*}$ have multiplicities, we factorise $\mathcal{F}=\mathcal{F}_{2}\circ\mathcal{F}_{1}$ as we did for $\mathcal{G}$ in eq.
 (43), so that we have the decompositions
@@
 $\operatorname{\operatorname{tr}}(\mu\mathcal{F}_{1}\otimes\mathrm{Id}(\sigma))$ $=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\Bigl(\sum_{\nu,\eta=1}^{d}|\nu\nu\rangle\!\langle\eta\eta|\Bigr)\Bigl(\sum_{\nu^{\prime},\mu^{\prime}=1}^{d}\mathcal{F}_{1}(|\nu^{\prime}\rangle\!\langle\eta^{\prime}|)\otimes|\nu^{\prime}\rangle\!\langle\eta^{\prime}|\Bigr))$
-$=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\sum_{\nu,\eta,\nu^{\prime},\eta^{\prime}=1}^{d}|\nu\nu\rangle\!\langle\eta\eta||\nu^{\prime}\nu^{\prime}\rangle\!\langle\eta^{\prime}\eta^{\prime}|)$
+$=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\sum_{\nu,\eta,\nu^{\prime},\eta^{\prime}=1}^{d}|\nu\nu\rangle\!\langle\eta\eta|\nu^{\prime}\nu^{\prime}\rangle\!\langle\eta^{\prime}\eta^{\prime}|)$
 $=\frac{1}{d^{2}}\bigl(\sum_{\eta,\nu^{\prime}=1}^{d}\langle\eta\eta|\nu^{\prime}\nu^{\prime}\rangle\bigr)\bigl(\sum_{\nu,\eta^{\prime}=1}^{d}\langle\eta^{\prime}\eta^{\prime}|\nu\nu\rangle\bigr)$
@@
 Let $\mathcal{F}$, $\mathcal{G}$ be two idempotent channels, $(k_{\nu})_{\nu\in\mathbb{N}}$, $(n_{\nu})_{\nu\in\mathbb{N}}$ be two integer sequences and $\operatorname{\varepsilon}>0$ satisfying the assumptions of Theorem 2.
-Let $p_{\min}\in\{1,+\infty\}$ such that
+Let $p_{\min}\in\{1,+\infty\}$ be such that
 $\min_{p\in\{1,+\infty\}}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}=\frac{\log(\|\lambda(\mathcal{G})\|_{p_{\min}})}{\log(\|\lambda(\mathcal{F})\|_{p_{\min}})}.$
@@
 Acknowledgements
-We thank Robert Salzmann for helpful discussions about [9] and [10], as well as about sets of approximate fixed point of channels.
+We thank Robert Salzmann for helpful discussions about [10] and [11], as well as about sets of approximate fixed points of channels.
+We thank Satvik Singh and Arturo Konderak for pointing out a mistake in Proposition 1 in the first version of the present article.
+We also thank Arturo Konderak for drawing our attention to [1].
 ID and OF acknowledge funding by the European Research Council (ERC Grant AlgoQIP, Agreement No.
@@
 $0\leq t^{2}Q+tV+W\leq t^{2}\delta\|x\|^{2}\mathds{1}+tV+\delta\|y\|^{2}\mathds{1}.$
-As $V=v+v^{*}$ is Hermitian, we have $V=UDU^{*}$ with $U$ a unitary matrix and $D$ a diagonal matrix with real coefficients. Conjugating the last inequality by $U$, we get:
+As $V=v+v^{*}$ is Hermitian, we have $V=UDU^{*}$ with $U$ a unitary matrix and $D$ a diagonal matrix with real coefficients. Conjugating the last inequality by $U$, we get
 $0\leq t^{2}\delta\|x\|^{2}\mathds{1}+tD+\delta\|y\|^{2}\mathds{1}.$
@@
 $\delta\|x\|^{2}t^{2}+D_{i}t+\delta\|y\|^{2},$
-has to be non-positive, that is
+has to be non-positive for all indices $i$, that is
 $D_{i}^{2}\leq 4\delta^{2}\|x\|^{2}\|y\|^{2}.$
@@
 […]
-$\|\mathcal{F}\|_{\text{cb}}=\sup_{n\in\mathbb{N}^{*}}\|\mathcal{F}\otimes\mathrm{Id}_{n}\|_{\infty\rightarrow\infty},$
+$\|\mathcal{F}\|_{\mathrm{cb}}=\sup_{n\in\mathbb{N}^{*}}\|\mathcal{F}\otimes\mathrm{Id}_{n}\|_{\infty\rightarrow\infty},$
 with $\|\cdot\|_{\infty\rightarrow\infty}$ the ‘infinity-to-infinity’ norm, i.e.
@@
 […]
-$\|\mathcal{F}\|_{\diamond}=\|\mathcal{F}^{*}\|_{\text{cb}}.$
+$\|\mathcal{F}\|_{\diamond}=\|\mathcal{F}^{*}\|_{\mathrm{cb}}.$
 Lemma 15 (Almost norm preservation).
 Let $\mathcal{F}$, $\mathcal{G}$, $\widetilde{\mathcal{G}}$, $\mathcal{E}$ and $\mathcal{D}$ be defined as in the statement of Proposition 8 and let
-$\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\text{cb}}.$
+$\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\mathrm{cb}}.$
 Then, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
@@
 $\overset{(c)}{=}\|\mathcal{F}^{*}(x)-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}(x)\|$
-$\leq\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{cb}\|x\|$
+$\leq\|\mathcal{F}^{*}-\mathcal{E}^{*}\mathcal{G}^{*}\mathcal{D}^{*}\|_{\mathrm{cb}}\|x\|$
 $\leq\delta\|x\|,$
@@
 where $(a)$ follows from the fact that as $x\in\operatorname{Rg}(\mathcal{F}^{*})$, we have
-$\|\Phi(x^{*}x)-x^{*}x\|=\|\Phi^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*}x)\|\leq\|\Phi^{*}-\mathcal{F}^{*}\|_{\text{cb}}\|x^{*}x\|\leq\delta\|x\|^{2}.$
+$\|\Phi(x^{*}x)-x^{*}x\|=\|\Phi^{*}(x^{*}x)-\mathcal{F}^{*}(x^{*}x)\|\leq\|\Phi^{*}-\mathcal{F}^{*}\|_{\mathrm{cb}}\|x^{*}x\|\leq\delta\|x\|^{2}.$
 By the same argument, we also have that $\|x-\Phi^{*}(x)\|\leq\delta\|x\|$ and $\|x^{*}-\Phi^{*}(x^{*})\|\leq\delta\|x\|$, hence $(b)$ follows.
@@
 Proof of Proposition 8.
-Let $\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}.$ By Lemmas 1415 and 17, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\max\{\delta,\frac{6d\delta}{\lambda_{\min,1}}\}$-inclusion (we trivially have that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is Hermitian preserving).
+Let $\delta\geq\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}$.
+Since we assumed that $\mathcal{F}$ and $\mathcal{G}$ are equal to their reduced channels, i.e.
+$\mathcal{F}=\widehat{\mathcal{F}}$ and $\mathcal{G}=\widehat{\mathcal{G}}$, $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$.
+By Lemmas 1516 and 18, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is a $\max\{\delta,\frac{6d\delta}{\lambda_{\min,1}}\}$-inclusion (we trivially have that $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ is Hermitian preserving).
 However, as
@@
 Proof of Theorem 11.
-Basically, we simply apply Proposition 7 to $\mathcal{F}^{\otimes k}:\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})$} and $\mathcal{G}^{\otimes n}:\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$.
+Basically, we simply apply Proposition 8 to $\widehat{\mathcal{F}}^{\otimes k}:\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes k})$} and $\widehat{\mathcal{G}}^{\otimes n}:\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$, with $\widehat{\mathcal{F}}$ and $\widehat{\mathcal{G}}$ the reduced channels respectively associated to $\mathcal{F}$ and $\mathcal{G}$.
 Just note that the dimension of $\mathcal{H}_{\mathcal{F}}^{\otimes k}$ is $d^{k}$, with $d=\dim(\mathcal{H}_{\mathcal{F}})$.
 Thus, by Proposition 8, if the hypotheses of Theorem 11 are met, for all $p\in[1,+\infty]$, we have:
-$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}\leq\|\lambda(\mathcal{G}^{\otimes n})\|_{p}.$
-This imply, by the multiplicativity of the $\ell_{p}$-norms of the shape vector with the tensor product, that
+$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}=\|\lambda(\widehat{\mathcal{F}}^{\otimes k})\|_{p}\leq\|\lambda(\widehat{\mathcal{G}}^{\otimes n})\|_{p}=\|\lambda(\mathcal{G}^{\otimes n})\|_{p}.$
+This implies, by the compatibility of the $\ell_{p}$-norms of the shape vector with the tensor product, that
 $\frac{k}{n}\leq\inf_{p\in[1,+\infty]}\frac{\log(\|\lambda(\mathcal{G})\|_{p})}{\log(\|\lambda(\mathcal{F})\|_{p})}=C(\mathcal{G}\mapsto\mathcal{F}),$